### v1

# Prioritizing Security Practice Adoption: Empirical Insights on Software Security Outcomes in the npm Ecosystem

Abstract.
Practitioners often struggle with the overwhelming number of security practices outlined in cybersecurity frameworks for risk mitigation. Given the limited budget, time, and resources, practitioners want to prioritize the adoption of security practices based on empirical evidence. The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices and security outcome metrics. The study investigated the relationship between security practice adoption and security outcomes. We selected the OpenSSF Scorecard metrics to automatically measure the adoption of security practices in npm GitHub repositories. We also explored security outcomes metrics, such as the number of open vulnerabilities (Vul_Count), mean time to remediate (MTTR) vulnerabilities in dependencies, and mean time to update (MTTU) dependencies. We conducted regression and causal analysis using 12 Scorecard metrics and their aggregated Scorecard score (computed by aggregating individual security practice scores) as predictors and Vul_Count, MTTR, and MTTU as target variables. Our initial dataset included 146K npm packages that have Scorecard, Vul_Count , and MTTU data. However, 85% lacked MTTR data because their dependencies either had no reported vulnerabilities or had vulnerabilities that remained unpatched. To address this, we extracted a 22K subset from the 146K dataset with MTTR data. Our findings show that higher aggregated Scorecard scores are associated with fewer Vul_Count and shorter MTTU, also supported by causal analysis. However, while the regression model suggests shorter MTTR, causal analysis indicates project characteristics likely influence MTTR direction. Segment analysis shows that larger, newer repositories with more contributors, dependencies, and downloads have shorter MTTR. Among individual security practices, Code Review, Maintained status, Pinned Dependencies, and Branch Protection show strong associations with security outcomes; the directionality of these associations varies across security outcomes. Additionally, MTTU follows trends similar to Vul_Count and MTTR, making it a potential proxy metric when vulnerability and MTTR data are unavailable.
1. Introduction
The package release frequency has increased by 1,466% over the last decade, with smaller applications now incorporating an average of 180 components (sonatype2024, ). The growing reliance on third-party dependencies in modern applications expands the attack surface, exposing applications to CVE (Common Vulnerabilities and Exposures catalog) exploits or malicious code injection, facilitating software supply chain (SSC) attacks (williams2024research, ). For example, CVEs have grown by 463% in the last 10 years, and 704,102 malicious packages have been proactively identified since 2019 (sonatype2024, ).
Practitioners are falling behind in managing vulnerabilities in dependencies due to growing dependency complexity and resource constraints, with mean time to remediate (MTTR) increasing from under 25 days in 2017 to over 400 days by 2024 (sonatype2024, ). SSC security standards have been proposed by industry and government agencies to help organizations adopt security practices that mitigate SSC security risks. Security practices are actions, procedures, techniques, or other measures that an organization can adopt to reduce the vulnerability of an information system (NISTSecurityControls, ). Examples of security practices include conducting code reviews before merging pull requests and using a static analyzer tool to detect vulnerabilities. In this paper, we used the term “security practice,” also synonymous with security control, safeguards, and countermeasures (NISTSecurityControls, ).
Many stakeholders have recognized SSC security concerns; most prominently, the Executive Order (EO) 14028 (EO_2021, ) mandates the adoption and self-attestation of security practices to improve SSC security. Similarly, the European Union’s Cyber Resilience Act (CRA), approved in March 2024, mandates common cybersecurity standards for digital products, including incident reporting and automatic MTTR updates (ENISA_CRA_2024, ). In 2025, the US EO 14144 (EO2025Cybersecurity, ) strengthens the efforts by requiring- “Within 240 days of the date of this order,….the Director of NIST, shall evaluate common cybersecurity practices and security outcomes that are commonly used or recommended across industry sectors.” Security outcomes are the indicators of the performance or non-performance of security practices prescribed for an information system (EO2025Cybersecurity, ).
While the SSC security standards (EO_2021, ; ENISA_CRA_2024, ; EO2025Cybersecurity, ; souppaya2022secure, ; Microsoft_framework, ; Williams2024PSSCRM, ) contain a long list of security practices that practitioners should adopt, practitioners are often overwhelmed by the number of practices recommended in these standards (zahan2023software, ). Given the limited budget, time, and resources (sammak2023developers, ; hamer2025closing, ), practitioners want to prioritize adopting security practices based on empirical evidence showing that adopting security practices mitigates security risk. The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices and security outcome metrics. Our research directly aligns with federal mandates EO 14144. However, instead of identifying commonly adopted security practices, we investigate the relationship between security practices and outcomes, providing empirical evidence to understand which practices are associated with improved security. We address the following research questions (RQs) and hypotheses:
  • RQ1: Is higher adoption of security practices associated with measurable improvements in security outcomes?
  • RQ2: Which security practices are most strongly associated with security outcomes?
To that end, the study investigated the npm ecosystem to understand the relationship between publicly available data on security practice adoption and security outcomes. We selected the OpenSSF Scorecard (Scorecard, ) tool for our large-scale analysis because the tool utilizes metrics that automatically measure the adoption of security practices in the GitHub repositories. The Scorecard (Scorecard, ) uses 18 predefined metrics to detect the adoption of security practices. For each metric, the Scorecard tool assigns a score of 0-10 based on pre-defined criteria. The tool also computes an aggregate security score as a weighted average of the 18 individual security metrics to understand the overall security of the repositories. For security outcomes, we selected the number of open vulnerabilities (Vul_Count), the mean time to remediate (MTTR), and the mean time to update (MTTU) metrics. To address RQ1, we developed a regression model to assess the relationship between a package’s aggregate security score and security outcome metrics. Additionally, we applied Propensity Score Matching (PSM), a statistical matching technique used in observational studies, to estimate causal effects and mitigate the potential confounding effects. For RQ2, we employed a regression model and feature importance-ranked techniques to identify which security practices have the strongest associations with security outcomes. To further validate these findings, we used PSM to estimate the effect of individual security practices on security outcomes, isolating their impact while accounting for differences in repository characteristics.
Our research builds on Zahan et al. (zahan2023software, ), which investigated whether OpenSSF Scorecard metrics predict vulnerabilities in the PyPI and npm ecosystems. Their study found a counterintuitive positive association between security practices and vulnerabilities and attributed the result to a small dataset (2,422 packages). However, the authors collected vulnerabilities at the package level while obtaining security practices from the GitHub repositories, which often contain additional dependencies and languages, leading to incomplete vulnerability data. We improved data collection by 1) incorporating vulnerabilities from dependencies and repositories instead of package level; 2) constructing a large dataset (146k npm packages); 3) including control variables (project age, size, etc.) to address external factors that might have an influence; and 4) introducing MTTR and MTTU as novel security outcome metrics to provide a more comprehensive evaluation. The study will have the following contributions:
  • MTTR and MTTU metrics to measure security outcomes.
  • A proposed model that can be used to explore the relationship between security practice and security outcome metrics.
  • Empirical evidence on how security practices impact security outcome of npm packages.
The paper is organized as follows: Section 2 discusses metrics used in our study, and Section 3 discusses our dataset construction steps. Section 4, Section 5 and Section 6 describes the data processing, methodology and result. We close with a discussion of our study (Section 7) and related work in Section 8.
2. Metrics Utilized in the Study
In this section, we discuss our metrics selection process, including security practices, security outcomes, and control variables.
2.1. Security Practice Metrics
In this study, we conducted a large-scale analysis to understand the adoption of security practices in OSS. We used the OpenSSF Scorecard tool (Scorecard, ), which evaluates security practices that can be measured automatically in GitHub. The tool detects security practice adoption, such as whether the package has implemented code review, repositories have enabled branch protection, or the repo is actively maintained. Table  1 contains the list of 18 security practices measured by Scorecard. Beyond industry adoption, the primary academic research on OpenSSF Scorecard has been conducted in study (zahan2023openssf, ; zahan2023software, ). The authors examine the adoption trends of security practices across ecosystems as measured by Scorecard metrics, analyze the distribution of Scorecard scores, and evaluate the efficacy of the Scorecard tool in assessing security practices. The author also verified whether the practices recommended by the Scorecard tool are also recommended in the standard security framework SSDF provided by NIST. The P-SSCRM SSC Framework (Williams2024PSSCRM, ) also included the OpenSSF Scorecard as one of the industry standards.
2.2. Security Outcome Metrics
Security outcomes are measured through quantitative characteristics, providing insights into the effectiveness of security practice adoption (scottmadden_outcome_metrics, ). In this study, we selected two temporal vulnerability metrics that capture both exposures (Vul_Count: number of open vulnerabilities) and responsiveness (MTTR: mean time to remediate vulnerable dependencies) and one temporal software metric (MTTU: mean time to update dependencies) as security outcome metrics.
  • Vul_Count: A measure quantifying the number of publicly disclosed or reported open security vulnerabilities within a project and its dependencies.
  • MTTR: A measure of the average aggregated time a package uses outdated and vulnerable direct dependencies in its lifetime (rahman2024characterizing, ). MTTR quantifies how long vulnerable dependencies remain outdated after a new version with a security fix has been released.
  • MTTU: A measure of the average aggregated time a package uses outdated direct dependencies in its lifetime (rahman2024characterizing, ). MTTU quantifies how long dependencies remain outdated after a new version has been released.
The number of vulnerabilities is widely used in academic research, industry, and policy as outcome metrics (EO2025Cybersecurity, ; CISASecureSoftwareAttestation, ; morrison2014mapping, ; sec_outcome, ; morrison2015security, ; zahan2023software, ). MTTR is most commonly used to measure the package’s security (MTTR_1, ; MTTR_2, ), measuring the code dependencies security (MTTR_3, ) in the industry. The OpenSSF Secure Supply Chain Consumption Framework (S2C2F) framework (Microsoft_framework, ) includes MTTR as a key maturity indicator, while Sonatype (sonatype-2019, ) has used MTTR to classify secure OSS projects. However, measuring MTTR poses a challenge since many package dependencies either have no reported vulnerabilities or remain unpatched, limiting the dataset size. To mitigate this limitation, we also consider MTTU to be a proxy metric for MTTR. Rahman et al. (rahman2024characterizing, ) found a moderate correlation between MTTR and MTTU, suggesting that MTTU can used as a proxy when MTTR data is unavailable. We consider MTTU as an alternative indicator of MTTR: 1) MTTU data was derivable for all components; and 2) We hypothesize developers staying up to date on dependencies will also stay up-to-date on security updates because security updates are a subset of regular updates. Hence, packages will exhibit similar MTTR and MTTU values. The hypothesis, data availability, and the observed correlation between MTTR and MTTU (rahman2024characterizing, ) motivated us to consider MTTU as an additional security outcome metric.
2.3. Control Variable Metrics
Control variables are properties that researchers hold constant for all observations in an experiment to understand the true relationships between the predictor and target variables (FrostControlVariables, ). In this research, we evaluated the impact of security practices on security outcomes while including other factors that may also influence the outcome. We included five control variables to measure project characteristics. GitHub contributions count controls for developer activity, while package downloads account for package popularity. Repository age (years) controls the project’s development stage, and size (KB) controls the complexity of maintaining a larger package. Direct dependency count controls a project’s reliance on and exposure to external dependencies.
Table 1. Scorecard Security Practices Metrics
Rows with bolded security practice names indicate excluded metrics, either due to high computational expense (HCE) or because more than 90% of repositories had a score of -1. A score of -1 indicates that Scorecard could not get conclusive evidence of implementing practices, or perhaps an internal error occurred due to a runtime error. A score between 0-10 represents the extent of adoption of the security practice.
Security Practices Description Risk Label -1 score 0-10 score
Dangerous-Workflow Does the project avoid dangerous coding patterns in GitHub Action workflows? Critical 73.2% 26.8%
Branch-Protection Does the project use Branch Protection? High ¡1% (7) 99.9%
Code-Review Does the project practice code review before code is merged? High ¡1.4% (2071) 98.6%
Binary-Artifacts Is the project free of checked-in binaries? High ¡1% (3) 99.9%
Dependency-Update-Tool Does the project use tools to help update its dependencies? High HCE HCE
Maintained Is the project at least 90 days old and maintained? High 0 100%
Signed-Releases Does the project cryptographically sign releases? High 97.7% 2.3%
Token-Permissions Does the project declare GitHub workflow tokens as read only? High 74.4% 26.8%
Vulnerabilities Does the project have unfixed vulnerabilities (including depedencies)? High 0 100%
Pinned-Dependencies Does the project declare and pin dependencies? Medium 72.2% 27.8%
Packaging Does the project build and publish official packages from CI/CD, e.g. GitHub Publishing ? Medium 97.4% 2.5%
Fuzzing Does the project use fuzzing tools, e.g. OSS-Fuzz, QuickCheck or fast-check? Medium ¡1% (15) 99.9%
Static Application Security Testing (SAST) Does the project use static code analysis tools, e.g. CodeQL, LGTM (deprecated), SonarCloud? Medium ¡1% (203) 99.9%
Security-Policy Does the project contain a security policy? Medium ¡1% (3) 99.9%
CI-Tests Does the project run tests in CI, e.g. GitHub Actions, Prow? Low HCE HCE
Contributors Does the project have contributors from at least two different organizations? Low HCE HCE
License Does the project declare a license? Low ¡1% (2) 99.9%
CII-Best-Practices Has the project earned an OpenSSF (formerly CII) Best Practices Badge at the passing, silver, or gold level? Low ¡1% (12) 99.9%
3. Dataset Construction
The primary source of our dataset are the npm packages and their dependency graph and GitHub repositories. In Section 3.1, we discuss different sources used to construct our dataset. The package inclusion and exclusion criteria are discussed in Section 3.2. Section  3.3 discusses our security practice metrics data collection process using the OpenSSF Scorecard tool. Section 3.4 and Section 3.5 cover our security outcome metrics and control variables.
3.1. Data Source
To collect data, we leveraged multiple sources. We selected the npm ecosystem for this study because npm is the largest open-source ecosystem, comprising 37 million package versions, with 74% of open-source codebases using JavaScript (ossra2024, ; sonatype2024, ). Scorecard analyzes packages with source repositories hosted on GitHub (openssf_scorecard_2023, ) and periodically evaluates popular open-source projects. Within Scorecard’s dataset of selected popular packages, 64% are npm packages. Additionally, our security outcome metrics require analyzing project dependencies, specifically counting vulnerabilities in dependencies and measuring dependency update patterns for MTTR and MTTU. To support our analysis, we collected the dependency graph of npm packages and mapped packages to their corresponding GitHub repositories using deps.dev (OSI, ). Deps.dev, developed and hosted by Google, provides comprehensive insights into OSS, including source code locations, dependency graphs, package metadata, licenses, releases, and vulnerabilities. Another platform used in our study is OSV (Open Source Vulnerabilities) (osv_adv, ). The OSV platform is an aggregator of vulnerabilities from different security advisories. Scorecard and Deps.dev depend on OSV to collect vulnerability information for projects and their dependencies. We also used the Libraries.io platform (LibrariesIO, ), which collects publicly available OSS information scraped from the internet, including GitHub.
3.2. Package Inclusion and Exclusion Criteria
To construct our dataset, we systematically filtered, cleaned, and processed npm packages based on dependency relationships, repository availability, and data availability. The dataset construction followed the following steps:
  • Initial Dataset: We began with a total of 2,119,044 unique npm packages extracted from the Deps.dev BigQuery dataset (depsdevBigQuery, ). This dataset contains the mapping between npm packages to GitHub repositories and captures all dependency edges within the npm.
  • Packages with GitHub Repositories: Of these, 1,241,732 packages were mapped to GitHub repositories. Since we must map each package to a Scorecard score via a GitHub repo, we exclude 877,312 packages that lack repository information, either due to hosting outside GitHub or missing repository URLs.
  • Packages with Dependencies and Dependents: Further filtering identified 1,159,891 packages with at least one declared dependency, ensuring relevance to our security outcome metrics. Additionally, we are primarily interested in packages that practitioners may decide to use; hence, we refined the dataset further to include only packages that had at least one dependent, reducing the dataset to 264,413 packages.
  • Ensuring Unique Repositories: Scorecard generates metric scores based on security practice adoption at the repository level rather than for individual packages. While we collected vulnerability data at the repository level (Section 3.4), MTTR and MTTU can only be measured at the package level. Consequently, if a repository contains multiple packages, all packages share the same Scorecard score and vulnerability data but may have different MTTR and MTTU values. Such inconsistency leads to duplicate rows with identical Scorecard metrics but differing security outcome metrics, which could mislead our machine learning model. To resolve this issue, we filtered the dataset to include only repositories that map to a single package. After applying this filtering step, we retained 172,999 unique repositories.
3.3. OpenSSF Scorecard Data
Once we identified our set of 172,999 unique npm GitHub repositories, we collected OpenSSF Scorecard scores for each repository associated with a package. Our primary data source was the OpenSSF Scorecard dataset in Google BigQuery (ScorecardBigQuery, ). However, Scorecard only evaluates a subset of repositories, specifically popular packages and those explicitly requested by practitioners for inclusion. From Google BigQuery, we extracted Scorecard data for 761,602 repositories. However, not all of these repositories met our inclusion criteria mentioned in Section 3.2. For example, some lacked dependencies or dependents, while others had multiple packages mapped to single repositories, which was outside the scope of our study. Hence, many repositories in our dataset were not included in the original Scorecard dataset. To address these gaps, we ran Scorecard locally for any repositories missing from the BigQuery dataset, using the same configuration applied to populate BigQuery. After merging data from Google BigQuery and our local runs, we obtained Scorecard scores for 146,951 out of 172,999 repositories. The remaining repositories did not generate a score using Scorecard. Upon manual verification of 100 repositories, we found that Scorecard could not produce scores primarily because the repositories were either private or had been removed from GitHub.
Metrics inclusion and exclusion In this study, we studied 13 out of 18 Scorecard metrics, excluding CI-Tests, Contributors, Dependency-Update-Tool, Signed Releases, and Packaging, due to computational constraints or a lack of sufficient data. The CI-Tests, Contributors, and Dependency-Update-Tool metrics are not included in the Google BigQuery dataset due to the computational resources required to run these checks, which would hinder periodic updates. The exclusion was confirmed by the OpenSSF team, who specified the configuration in the project’s source code  [^1] https://github.com/ossf/scorecard/blob/76a04bfe40c25746b2633127270a9219a98a37ad/cron/config/config.yaml#L48. Scorecard generates a -1 score when no conclusive evidence of a practice’s adoption exists, either due to insufficient repository data or internal runtime errors. Signed Releases and Packaging were excluded, as over 90% of the repositories received a -1 score (Table 1). Prior research (zahan2023openssf, ) found that practitioners primarily used package registries for releases rather than GitHub, and no improvements in Scorecard’s methodology addressing this limitation were observed. Thus, we removed five metrics from the analysis.
Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions metrics have less than 75% repositories with a -1 score (Table 1). Prior research by Zahan et al. (zahan2023openssf, ) highlighted limitations in Scorecard’s tool, particularly for these metrics that are used to assess security configuration in GitHub workflows and dependency management practice (Table 1). Previously, Scorecard assigned a maximum score of 10 for these metrics, even when no GitHub workflows or dependency metadata were present, which could mislead practitioners about positive security practice adoption. Our manual analysis of 50 repositories per metric confirms that Scorecard has improved the detection of these practices. Scorecard assigns a -1 score when repositories do not have any GitHub workflows, token information, or dependency metadata. A score of 0 now signifies that the practice exists with security misconfiguration, such as excessive GitHub token permissions, dangerous workflow patterns, or all dependencies are unpinned. We included these three metrics in our study because a -1 score represents non-adoption, while a score of 0 indicates security misconfigurations.
3.4. Security Outcome Metrics Data
We collected our security outcome metrics for 146,951 npm packages with Scorecard data. Here, we describe our outcome metrics data collection process.
Vul_Count metric: The Scorecard vulnerability metrics measure whether the project has open, unfixed vulnerabilities in its codebase or its dependencies using the OSV (osv_adv, ) service. Initially, we used Deps.dev BigQuery (depsdevBigQuery, ) and the Scorecard tool to collect the total vulnerability count. Deps.dev (OSI, ) includes Scorecard metrics for OSS. In deps.dev, we were able to map vulnerability data for 76,651 packages, and 70,300 remaining data were collected using Scorecard. The scorecard normalizes the vulnerability score, where a score of 10 indicates no vulnerabilities, 9 corresponds to one vulnerability, and 0 represents more than nine vulnerabilities. To obtain the exact vulnerability count, we first denormalized the score. For cases where the vulnerability score was 0 (indicating more than nine vulnerabilities), we ran the vulnerability metric separately using Scorecard to retrieve the raw, unnormalized data.
MTTR and MTTU metric: In this study, we adopt the algorithms proposed by Rahman et al. (rahman2024characterizing, ) for computing MTTR and MTTU. Following their methodology, we contacted Google deps.dev team and considered deps.dev as our primary data source to collect comprehensive package versions and dependency versions and their release information for our dataset. Additionally, we retrieved CVE data for selected packages and their direct dependencies from OSV (osv_adv, ). Note that while MTTU captures all dependency updates, MTTR specifically accounts for updates addressing vulnerable dependencies. While we successfully collected Vul_Count and MTTU data for 146,951 packages, MTTR data was only available for 22,083 packages (15% of total dataset). The limitation arises because MTTR penalizes packages that did not update their dependencies despite a known vulnerability being fixed in an upstream version.
3.5. Control Variable Metrics Data
We collected control variable data from different sources. For example, GitHub contribution count and package size are collected from Libraries.io (LibrariesIO, ). Then we got the number of downloads of a package in the past 12 months, which we collected from the public npm API [^2] https://api.npmjs.org/downloads/point/{period}[/{package}]. We calculated the repository age (in years) using the GitHub REST API (GitHubAPI2024, ) by retrieving the repository creation date and computing the difference between the current date and the creation date.
4. Data Preprocess
Before constructing machine learning (ML) modeling to answer RQ1, we performed data preprocessing steps to handle missing values, ensure data integrity, and assess statistical properties. Section 4.1 covers data imputation techniques to manage missing values. Section 4.2 discusses correlation and multicollinearity in our dataset. Figure 1 highlights the overview of our research workflow.
Figure 1. Research Workflow
4.1. Data Imputation
Data imputation is the process of replacing missing values to maintain data integrity, prevent biases in ML models, and ensure statistical validity (schafer2002missing, ). Imputation techniques help mitigate data loss, improve model accuracy, and prevent misleading conclusions caused by missing data. Depending on the nature of the missing data, multiple imputation techniques exist, such as mean/median imputation, categorical encoding, and predictive modeling-based imputations. In our dataset, Branch-Protection, Code-Review, Binary-Artifacts, Fuzzing, SAST, Security-Policy, License, and CII-Best-Practices each had less than 1% of packages with a -1 score. Table 1 shows the exact counts of packages with -1 for these metrics. On the contrary, Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions had greater than 72% of repositories with a -1 score (Table 1). Our manual analysis found that Scorecard assigned -1 in two cases: (1) when there was no conclusive evidence of security practice adoption, and (2) due to internal runtime errors (section 3.3). Since not all metrics within a repository had a -1 score, removing entire repositories would result in unnecessary data loss. To handle missing data effectively, we differentiated between cases where -1 indicated non-adoption (e.g., missing GitHub workflows, token permissions, or dependency management practices) and cases where -1 resulted from internal runtime errors. For metrics where -1 was due to internal errors, we applied median imputation to maintain data completeness while minimizing distortion. For Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions, where -1 indicated non-adoption rather than missing data, we applied categorical encoding imputation to preserve the distinction. Note that a score of 0 reflects detected security misconfigurations.
Figure 2. Spearman Correlation between Predictors
4.2. Correlation Analysis
To evaluate dataset normality, we applied the Henze-Zirkler multivariate normality test(henze1990class, ) separately to the RQ1 and RQ2 datasets  (Figure 1). The RQ1 dataset includes aggregated Scorecard score, Control Variables, Vul_Count, MTTR, and MTTU, while RQ2 leverage 12 Scorecard Metrics instead of aggregated Scorecard score. Both datasets failed the normality test (p ¡ 0.05), indicating deviations from a normal distribution. To further assess skewness, we flagged features with —skewness— ¿ 1 for potential transformation (hatem2022normality, ). Based on this, we applied log(x + 1) transformation to highly skewed variables (e.g., GitHub contributors, repository size, DependencyCount, downloads, MTTU, and Vul_Count) to mitigate skewness. Post-transformation, Q-Q plots, histograms, and statistical tests confirmed improved distribution symmetry. Since residual normality is important for regression analysis, we conducted D’Agostino-Pearson tests (keskin2006comparison, ) on regression residuals. Results showed that residuals remained non-normally distributed (p ¡ 0.05).
Given the non-normal distribution of our dataset, we proceeded with the Spearman correlation (zar2005spearman, ), a nonparametric method suitable for capturing monotonic relationships in non-normal data. We conducted correlation analysis for predictor variables (security practice and control variables) and target variables (security outcome metrics). Figure  2 presents the Spearman correlation heatmap for predictor variables, revealing weak to moderate correlations. The highest correlation is observed between contributors and code review (0.61), suggesting that repositories with more contributors are more likely to implement code review practices. The correlation between target variables reveals a strong positive correlation (0.7) between MTTR and MTTU, aligning with our expectation of potential overlap (rahman2024characterizing, ). In contrast, the Vul_Count metric exhibits a weak negative correlation with MTTR (-0.32) and MTTU (-0.04). Hence, we confirmed that MTTR and MTTU cannot be used as target variables in the same model due to their strong correlation since including both would introduce redundancy, bias, and multicollinearity. Additionally, we computed the Variance Inflation Factor (VIF) (o2007caution, ) for all predictor variables to confirm multicollinearity. For VIF computation, we found only “downloads” metrics had a high VIF score (15). According to study (o2007caution, ), VIF thresholds should be interpreted in context rather than strictly applied as exclusion criteria. Since the ”download” metric serves as a control variable in our study and does not exhibit strong correlations (¿0.8) with any other metrics, the high VIF alone is not sufficient grounds for removal (o2007caution, ).
5. RQ1: Overall security posture and its security outcomes
In this section, we discuss our RQ1: Is higher adoption of security practices associated with measurable improvements in security outcomes? Section 5.1 discusses our methodology to answer our RQ1, and Section 5.2 discusses our result.
Table 2. Summary Statistics of Security Outcomes
Statistic Vul_Count MTTR MTTU
Count 146,951 22,543 146,951
Mean 16.4 1,084.7 634.3
Std Dev 31.3 897.3 877.4
Min 0 1 0
Median 1.0 940.0 222.2
Max 637.0 4,709.0 4,869.6
5.1. RQ1: Methodology
For our RQ1, we evaluate the relationship between npm package repositories’ aggregated Scorecard score and control variables as predictors and security outcome metrics as the target variable (Figure 1). Table 2 presents the summary statistics of security outcomes. To visualize the relationship between aggregated Scorecard scores and security outcomes, we first binned repositories by aggregated Scorecard scores and analyzed mean security outcomes within each bin (Figure 3). Next, we applied regression analysis to quantify the association of the aggregated Scorecard score on outcomes. Finally, we conducted a causal analysis to assess whether the observed relationships reflect a causal effect.
5.1.1. Regression Model:
We selected Generalized Linear Models (GLMs) (nelder1972generalized, ), given the non-normal distribution of our dataset and our target variables are count data. GLMs are specifically designed for count data and provide interpretable coefficients. Coefficients quantify the expected change in the target variable for a unit increase in a predictor. To identify the best-fitting GLM, we systematically evaluated the dataset characteristics using a data-driven approach. Our selection process was guided by two factors: the Overdispersion Test (variance greater than the mean) and the zero-inflation test(a large proportion of zero values in target variables). We first started with the Poisson Regression (hausman1984econometric, ) as our baseline model and found that the dataset contains overdispersion for Vul_Count , MTTR, and MTTU, and zero inflation exists for Vul_Count only. We selected Negative Binomial (NB) Regression (hilbe2011negative, ) for MTTR and MTTU to address overdispersion and Zero-Inflated Negative Binomial (ZINB) (ridout2001score, ) to handle both overdispersion and zero inflation for Vul_Count. Since GLMs inherently apply a log link function instead of log transformation, we used the raw target variables to maintain the interpretability of the coefficients.
Figure 3. Security Outcome Trends and npm Repository Distribution Across Aggregated Scorecard Score Bins
5.1.2. Causal Analysis:
While regression models can identify associations between “aggregated Scorecard score” and security outcomes, they do not establish causality due to the potential influence of confounding factors. For example, the predictors aggregated Scorecard score and control are correlated, which could bias the estimation of the true impact of security practices on security outcomes. To address this, we used Propensity Score Matching (PSM) (austin2009balance, ), a statistical technique that simulates a randomized experiment by pairing repositories with similar characteristics. PSM matched the treated units (repositories with aggregated Scorecard scores above the median) with control units (below the median) based on covariates. Covariates are control variables that could influence both aggregated Scorecard scores and security outcomes. Using logistic regression, we estimated each repository’s likelihood of having a high aggregated Scorecard score. We then applied nearest-neighbor matching with a caliper adjustment(0.1) to pair treated repositories with control repositories based on similar covariates. The similarity of matching was determined using propensity scores, which represent the likelihood that a repository belongs to the treated group based on the repository characteristics. To validate the matching process, we used standardized mean differences (SMD) to assess balance across covariates. Finally, we calculated the Average Treatment Effect (ATE) using an inverse probability-weighted (IPW) estimate (ATE weighted) that adjusts for any residual differences. Computing the Average Treatment Effect (ATE), we estimated how much a package’s security outcome would change if it had a higher aggregated Scorecard score.
5.2. RQ1: Result
Figure 3 shows a sharp decline in security outcomes (in Vul_Count, MTTR, and MTTU) as the aggregated Scorecard score increases. The highest bin (7.3–9.1) reflects the top-scoring repositories, as no project achieved a perfect score. However, the plot only reflects an overall trend and does not account for confounding factors and associations. To formally assess the relationship, we use regression analysis and propensity score matching (PSM). Table 3 demonstrates that a higher aggregated Scorecard score is associated with improved security outcomes (negative coefficient). In the ZINB model, a coefficient of –1.4 for Vul_Count implies that a unit increase in the aggregated Scorecard score corresponds to a 1.4 count decrease in the expected number of vulnerabilities. The finding is also supported by negative ATE estimates (–1.9) in the PSM analysis. For MTTU, both the NB model (coefficient –0.2) and PSM estimates (–0.5) indicate that repositories with higher aggregated Scorecard scores update dependencies more rapidly. By contrast, the NB model for MTTR shows a downward trend (Figure 3) and a negative coefficient (–0.1). However, the PSM analysis yields a positive overall ATE (74.7), indicating that, when controlling for control variables, repositories with higher aggregated Scorecard scores tend to have longer update delays for vulnerable components.
Table 3. GLMs and PSM analysis between Aggregated Scorecard Score and Security Outcomes
Dep. Var Model Coef. p-value ATE(Weighted)
Vul_Count ZINB -1.4 ¡0.001 -1.9
MTTR NB -0.1 ¡0.001 74.7
MTTU NB -0.2 ¡0.001 -0.5
We conducted a segment-based analysis for MTTR to investigate further the discrepancy by stratifying the dataset into subgroups based on repository characteristics using a median split. The segment analysis revealed that projects in the upper groups for size, contributors, dependency count, and downloads and the lower group for age tend to exhibit negative ATEs (indicating faster MTTR). In contrast, repositories in the lower groups for size, contributors, dependency count, and downloads and in the upper age group show positive ATEs (indicating slower MTTR). Overall, the segment analysis reveals that the effectiveness of security practices adoption in reducing MTTR is context-dependent. Larger, newer repositories with more contributors, dependencies, and downloads benefit from adopting security practices, resulting in faster MTTR. In contrast, smaller but older repositories with fewer contributors, downloads, and dependencies may experience increased overhead from adopting security practices, leading to longer MTTR.
Summary: Our analysis shows that higher aggregated Scorecard scores are associated with fewer vulnerabilities and shorter MTTU. However, while regression models indicate shorter MTTR, causal analysis reveals that repositories with higher aggregated Scorecard scores may take longer to update vulnerable components since project characteristics also influence the relationship.
Table 4. Distribution of Security Practices Adoption
Metric % -1 % 0 % 1-10
Branch Protection 0.005% 85.9% 14%
Code Review 1.4% 74.6% 24%
License 0 31% 69%
Maintained 0.001% 86.7% 13.3%
Pinned Dependencies 72.2% 0 27.9%
Table 5. R2, Security practice Importance Ranked and PSM Insights
Dep. Var R2 for RF Ranked Security Practices PSM Effect (ATE weighted direction)
Vul_Count 0.38 Code Review, Maintained, Branch Protection, License, Pinned-Dependencies Code Review (+), Others (-)
MTTR 0.42 Maintained, Pinned-Dependencies, Branch Protection, Code Review, License All (+)
MTTU 0.2 Maintained, Code Review, Pinned-Dependencies, Branch Protection, Security Policy Code Review (+), Branch Protection(+), Others (-)
6. RQ2: Associations between security practices and outcomes
In this section, we discussed our RQ2: Which security practices are most strongly associated with security outcomes? Section 6.1 discusses our methodology to answer our RQ2, and Section 6.2 discusses our result.
6.1. RQ2: Methodology
For RQ2, we investigate the association between individual security practices and security outcomes. Unlike RQ1, where we analyzed aggregated Scorecard scores as a predictor, for RQ2, we analyzed twelve distinct security practices as predictors. The large number of predictors, non-normal distribution, and non-linear relationships among variables (section 4) introduce additional complexity to the dataset. We selected a tree-based regression model as these models are well-fitted for datasets with complex, non-linear relationships and high-dimensional feature interactions. We also conducted the feature importance ranked analysis to identify the security practices that most strongly contribute to predicting security outcomes. We performed causal analysis to explore potential causal relationships since regression and feature importance identify associations but do not establish causality.
6.1.1. Regression Model:
Among tree-based models, we chose Random Forest (RF) (dewi2019random, ) due to its ensemble learning approach, which aggregates multiple decision trees to improve model performance and reduce overfitting. Unlike single decision trees, which can be highly sensitive to variations in training data, RF mitigates the issue by averaging predictions across multiple trees. We also utilized Gini Importance (Mean Decrease in Impurity) as a feature importance technique to quantify each security practice’s contribution toward reducing prediction error. Gini Importance measures a predictor’s contribution by assessing how often the predictor is used for splitting the data in the decision trees of the RF model, highlighting the practice’s contribution in accurately predicting security outcomes. We first preprocessed our data, converting categorical security practices into appropriate numerical or binary formats (section 4.1). We then split the dataset into training (70%) and testing (30%) subsets. The model was trained on the training dataset using default hyperparameters. Five-fold cross-validation was applied to estimate and reduce bias, calculating mean squared error (MSE) and R² scores. The final model’s performance and feature importance scores were validated on the reserved testing set (Table 5).
6.1.2. Causal Analysis
We followed the methodology outlined in Section 5.1.2 for our causal analysis. We conducted the PSM analysis only on the top 5 security practices ranked by the Gini feature importance measure. The top 5 was selected because, after that, the feature importance value was almost near 0, indicating those practices did not have a strong association with security outcomes. The aggregated Scorecard score is computed as a weighted average of 18 individual security practices; only 25 repositories resulted in a score of zero. On the contrary, individual security practices exhibit a higher rate of non-adoption or security misconfiguration (score -1, score 0), as shown in Table 4 (top 5 based on feature ranked). The disparity occurs because aggregation combines multiple practices, ensuring a nonzero score even when some practices are not adopted. In contrast, an individual practice remains at 0 or -1 if the practice is not implemented. As a result, individual security practices had a higher proportion of repositories with a score of 0 or -1, leading to a highly imbalanced dataset for PSM analysis. To mitigate imbalance and ensure balanced treatment and control groups, we applied a sampling approach for the top 5 security practices. We first included all repositories with scores between 1 and 10, then randomly sampled an equal number of repositories with a practice score of zero to create a representative baseline. We then created a binary treatment variable by splitting the combined dataset based on an adaptive threshold (below and above the median score). However, for Maintained and License metrics, the median score was equal to the maximum observed score (indicating minimal variance), and we adjusted the threshold to the midpoint between the minimum and maximum scores to ensure separation between high and low adoption groups. After constructing balanced groups, we followed the technique discussed in Section 5.1.2, where we used nearest-neighbor matching with a caliper adjustment and computed the ATE score (ATE weighted).
6.2. RQ2: Result
Our analysis employed RF regression, complemented by feature importance analysis and Propensity Score Matching (PSM), to identify security practices most strongly associated with security outcomes. Table  5 summarizes our findings, illustrating the variance explained (R²) by security practices and each practice’s directional effects based on the PSM analysis’s ATE-weighted values. The R² value indicates that security practices collectively explain a moderate portion of the variance in Vul_Count (R² = 0.38) and MTTR (R² = 0.41), and a relatively smaller proportion for MTTU (R² = 0.2). Notably, the lower R² for MTTU implies that additional external factors or project-specific characteristics beyond those captured in our current model likely contribute to MTTU. From the feature importance analysis, Code Review, Maintained, Pinned Dependencies, and Branch Protection practices consistently rank highly across all three outcome metrics. License considerations showed relevance for both Vul_Count and MTTR, whereas Security Policy appears influential for MTTU, highlighting the variability of practice impacts across distinct outcome measures.
In the PSM analysis of individual security practices, we did not consistently observe a negative (i.e., improving) effect across all security outcomes. Most security practices show a negative ATE for Vul_Count and MTTU (Table  5), implying fewer vulnerabilities and faster update frequency. However, Code Review consistently exhibits a positive relationship with all three security outcomes, suggesting that the practice aids in detecting more vulnerabilities. As emphasized in CISA’s Secure by Design guidance (cisa2023securebydesign, ), a rise in reported vulnerabilities can reflect a healthy security posture due to code review and testing by the community. For MTTR, all practices exhibit a positive ATE, indicating longer remediation times for repositories adopting these practices. Figure 3 shows that 90% npm repositories fall into low-security score bins (0.0–3.6), where MTTR is high (1200 days), suggesting these projects do not prioritize dependency security or vulnerability remediation. Table 4 further highlights that many npm packages lack security practices adoption, reinforcing this trend. While higher aggregated security scores and individual practices are associated with fewer Vul_Count and shorter MTTU, MTTR remains positive for both cases, suggesting MTTR is influenced by factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity. Unlike RQ1, where segmentation revealed how repository characteristics impact MTTR, we could not perform a similar analysis in RQ2 due to data imbalance (Section 6.1). The limited sample size in the PSM analysis further constrained our ability to isolate project-specific influences on MTTR. Therefore, we cannot conclusively confirm whether the observed increase in MTTR consistently indicates slower remediation or whether the effect might vary depending on project-specific characteristics, maintainability, or dependency complexity. Overall, Vul_Count , MTTR, and MTTU exhibit similar trends as security outcomes metrics. However, MTTR is more complex and may have an influence on external factors such as project maintainability, resources, and dependency complexity that need further research.
Summary: Our analysis shows that Code Review, Maintained, Pinned Dependencies, and Branch Protection are strongly associated with security outcomes. While individual security practice is associated with fewer vulnerabilities and faster dependency updates, MTTR remains high, suggesting that factors beyond security practice adoption, such as project maintainability, resources, and dependency complexity, influence MTTR.
7. Discussion
The section discusses the novelty of our study (section 7.1) and the implications of our findings (section 7.2) and explores the threat to the validity (section 7.3) of our study.
7.1. Novelity of our study
Prior studies (zahan2023software, ; sonatype_8, ) have utilized OpenSSF Scorecard metrics alongside vulnerability count data to understand the relationship between security practices and security outcome metrics. Building on this foundation, we expand security outcome metrics to include MTTR and MTTU. Our findings showed that Vul_Count, MTTR, and MTTU are associated with security practices. However, MTTR is more complex and may have an influence on external factors such as project maintainability, resources, and dependency complexity.
Additionally, our study further advances existing research by addressing methodological limitations in prior work. Zahan et al. (zahan2023software, ) constructed a small dataset of 2,422 PyPI and npm packages and analyzed the relationship between OpenSSF Scorecard metrics and vulnerability counts. The authors’ findings revealed a positive association between security practice adoption and vulnerability counts, alongside a low explanatory power ($R^{2}$ 12%). The authors attributed these results to dataset size limitations, lack of vulnerable data, and the absence of control variables related to project characteristics. In contrast, our study leverages a larger and more representative dataset, encompassing approximately 146,951 npm package repositories, which mitigates biases introduced by small sample sizes and enhances generalizability. Additionally, we incorporate control variables, including project age, size, contributors, and dependency count and downloads, to account for potential confounding factors influencing security outcomes. Furthermore, Zahan et al.’s dataset had a temporal mismatch between security practice adoption and vulnerability data. The authors collected security practice scores at a single point in time but included all historically reported vulnerabilities throughout a package’s lifetime, potentially distorting the actual relationship between security practices and security outcomes. Our methodology addresses this inconsistency by aligning security practice adoption with vulnerabilities that actively existed at the time of our study. By ensuring a concurrent snapshot of security practices and vulnerabilities, our approach provides the temporally aligned evaluation of security practice effectiveness.
Additionally, we addressed the issue of data consistency in study (zahan2023software, ), stemming from repositories containing multiple packages, each potentially contributing different Vul_Count  (Section  3.2), despite sharing identical Scorecard scores. Moreover, vulnerabilities in prior studies (zahan2023software, ) did not include vulnerabilities from dependencies or entire repositories but rather only individual packages, introducing further inconsistencies. Such discrepancies result in duplicate predictors with varying outcome values, potentially biasing predictive models. To mitigate these issues, we filtered our dataset to include only repositories linked uniquely to a single package. We considered vulnerabilities explicitly associated with the repository itself and those inherited from dependencies.
In another study, Sonatype (sonatype_8, ) employed a random forest classification model to classify repositories as vulnerable or non-vulnerable using Security Scorecard metrics and project attributes. However, a binary classification approach is inherently limited in capturing relationships between security practices and security outcomes. The classification does not effectively measure the magnitude or direction of associations, nor does it indicate the relative strength or complexity of individual predictors. In contrast, our study utilizes regression and causal analysis, methods specifically designed to explore and quantify associations, thus providing insights into how variations in security practices systematically relate to security outcomes.
Summary: Overall, our methodological refinements, including extensive dataset size, the inclusion of project characteristics, causal analysis, targeted dataset inclusion strategies, and the introduction of MTTR and MTTU as additional novel security outcome metrics, offer a more precise and contextually generalizable evaluation of the relationship between security practices and outcomes.
7.2. Implications of our study
Our study provides actionable insights for practitioners, policymakers, and researchers by establishing measurable security outcome metrics to assess the effectiveness of security practices. For practitioners, our findings highlight that Code Review, Maintained, Pinned Dependencies, and Branch Protection are strongly associated with security outcomes. Organizations can prioritize these practices to improve their project security. By periodically tracking security outcome metrics, practitioners can evaluate whether their security interventions are effectively reducing vulnerabilities and improving remediation efficiency. Our research supports policymakers by providing empirical evidence on the effectiveness of security practices, informing data-driven policy decisions, and enhancing regulatory frameworks. For researchers, our study underscores the importance of incorporating longitudinal data, expanded covariate sets, and more granular temporal analyses to understand better the dynamic interactions between security practices, external factors, and security outcomes. Additionally, future research should explore new security outcome metrics beyond vulnerability counts and remediation times to develop a more comprehensive framework for measuring security improvements in software development.
7.3. Threats to validity
Our regression and causal analyses provide evidence supporting associations between aggregated Scorecard scores, security practices, and security outcomes within our dataset. However, several inherent biases in our dataset must be acknowledged. First, our analyses are based on snapshot data capturing repository Scorecard metrics score and characteristics at a single point in time, lacking temporal granularity. Without longitudinal or time-series analysis, uncertainty remains about whether changes in security practices precede or directly influence observed security outcomes, thereby limiting generalizability beyond our dataset. Second, selection bias may be present, as repositories that adopt higher security practices could systematically differ from those with lower adoption in unmeasured aspects such as organizational resources, developer expertise, or security culture. The unobserved factors may confound the relationships we examine, leading to potential over or underestimations of the true impact of security practices. Third, although PSM attempts to mitigate confounding factors by simulating randomized experiments, PSM inherently relies on observed covariates. Therefore, unobserved or missed variables such as organizational policies, economic incentives, developer motivations, or project complexity could influence both security practices and security outcomes. Additionally, the consistently observed positive ATE for MTTR suggests that security practices may introduce operational overhead, and the relationship is context-dependent, where repositories with project characteristics may experience varying impacts.
Apart from that, our dataset exhibits other imbalances, as most repositories have security practice scores of either 0 or -1, reflecting the limited adoption of these practices across the npm ecosystem. The real-world imbalance, characterized by few repositories adopting security practices, may introduce bias into our models, potentially skewing predictions or causal estimations. However, such imbalance accurately mirrors the current state of practice adoption within the npm, highlighting a widespread lack of security practice adoption among npm repositories. Hence, the imbalance reflects the actual state of adoption rather than a methodological flaw, highlighting the need for targeted interventions to increase security practice adoption among npm repositories. Finally, while our filtering approach improves data consistency by retaining only repositories uniquely linked to a single npm package and aligning vulnerability data at the repository level and MTTR and MTTU data at the package level, some inconsistencies may persist. For example, repositories that maintain multiple packages from different ecosystems or use different languages may still introduce variations in MTTR and MTTU value, potentially affecting model accuracy.
Summary: Overall, while our study provides insights into the relationships between security practices and security outcomes, the limitations highlight the need for further research. Future studies incorporating longitudinal data, expanded covariate sets, and more granular temporal analyses will be essential to better capture the dynamic interactions between security practices, external factors, and security outcomes.
8. Related Works
This section highlights prior related studies, existing security guidelines, frameworks, and security outcome metrics. Security Practice Frameworks: Executive Order (EO) 14028 emphasized adopting and self-attesting security practices to improve software security. To align with the EO, government agencies and industry experts have introduced various software supply chain security frameworks to enhance software supply chain security. These include the Software Component Verification Standard (SCVS) by OWASP (scvs, ), the Building Security In Maturity Model (BSIMM) (BSIMM, ), the CNCF Technical Advisory Group (TAG) recommendations (CNCF_TAG, ), Supply Chain Levels for Software Artifacts (SLSA) (SLSA, ), NIST Secure Software Development Framework (SSDF) (souppaya2022secure, ), the Open Source Software (OSS) Secure Supply Chain (SSC) Framework by Microsoft (Microsoft_framework, ), NIST’s Cybersecurity Supply Chain Risk Management Practices (NIST800-161, ), the CISA Secure Software Development Attestation Form (CISASecureSoftwareAttestation, ) and OpenSSF Scorecard (Scorecard, ). The Proactive Software Supply Chain Risk Management (P-SSCRM) (Williams2024PSSCRM, ) released in 2024, results from the study of real-world software supply chain risk management initiatives and the union of the 73 security practices in ten government and industry documents (standards and frameworks) mentioned above. In 2025, EO 14144 (EO2025Cybersecurity, ) was issued, emphasizing the need for NIST to evaluate common cyber practices and security outcomes across industry sectors.
Security Outcome Metrics: Various metrics can assess the effectiveness of security practices by tracking security performance over time, generally categorized as threat-oriented, asset/impact-oriented, and vulnerability-oriented (NIST800-55, ; NIST800-30, ). While threat- and asset/impact-oriented metrics offer valuable insights, they are difficult to quantify at scale due to reliance on real-time threat intelligence and complex impact assessments. In contrast, vulnerability-oriented metrics are more accessible and relevant to software security. Among them, severity-based metrics (e.g., CVSS) face challenges such as scoring inconsistencies and reliance on exploitability data (pendleton2016survey, ). Temporal vulnerability metrics, such as vulnerability count, time to detect, time to repair and time to remediate, are widely used to evaluate security (morrison2014mapping, ; sec_outcome, ; CISA-2023, ; morrison2018mapping, ). Morrison et al. (morrison2018mapping, ) conducted a systematic mapping study on software security metrics to evaluate the security properties of software and found that the most cited and most used metric is the vulnerability count (vulnerability metric). Davidson (davidson2009good, ) studied the time to close bug/vulnerability’ (a form of mean time to repair) as a security metric and measured its effectiveness. Beres et al. (beres2009using, ) studied the window of exposure’ (vulnerability metric), which is the length of interval of time needed by the security team to remediate a vulnerability once it has been disclosed. However, the scarcity of reported vulnerabilities, compounded by silently fixed issues, limits their reliability in assessing security outcomes (alhazmi2007measuring, ; dunlap2023finding, ). To address these limitations, prior work suggests alternative metrics, such as fault prediction models (shin2013can, ). Additionally, in software reliability, MTTR and MTTU have long been used to measure system resilience and maintenance efficiency (gokhale1999time, ; calvo2023applying, ). While MTTR is often employed in security contexts (MTTR_1, ; MTTR_2, ; MTTR_3, ), its reliance on reported vulnerabilities restricts its applicability. MTTU, by focusing on dependency updates, offers broader coverage but remains constrained by proprietary measurement methods (cogo2019empirical, ).
9. Conclusion
In this study, we investigated how aggregate Scorecard score and 12 Scorecard security practices metrics impact security outcome metrics. Our regression model and causal analysis reveal that higher aggregated Scorecard scores are associated with fewer vulnerabilities and shorter MTTU. However, while regression models indicate shorter MTTR, causal analysis suggests that factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity, influence MTTR. Feature ranking highlights Code Review, Maintained status, Pinned Dependencies, and Branch Protection as top security practices associated with security outcomes, though their associations with security outcomes vary in direction. Additionally, MTTU exhibits trends similar to vulnerability count and MTTR, making it a potential proxy metric when vulnerability and MTTR data are unavailable. Our research aligns with federal cybersecurity initiatives (EO 14144), by providing empirical evidence to guide security policy and industry practices. Future work should incorporate longitudinal data and additional security outcome metrics to refine further our understanding of how security practices impact software security over time.

### v2

@@
-Prioritizing Security Practice Adoption: Empirical Insights on Software Security Outcomes in the npm Ecosystem
+Assumptions to Evidence: Evaluating Security Practices Adoption and Their Impact on Outcomes in the npm Ecosystem
 Abstract.
@@
 Given the limited budget, time, and resources, practitioners want to prioritize the adoption of security practices based on empirical evidence.
-The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices and security outcome metrics.
-The study investigated the relationship between security practice adoption and security outcomes.
+The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices adoption and security outcome metrics.
+To do this, we analyzed the adoption of security practices and their impact on security outcome metrics across 145K npm packages.
 We selected the OpenSSF Scorecard metrics to automatically measure the adoption of security practices in npm GitHub repositories.
-We also explored security outcomes metrics, such as the number of open vulnerabilities (Vul_Count), mean time to remediate (MTTR) vulnerabilities in dependencies, and mean time to update (MTTU) dependencies.
-We conducted regression and causal analysis using 12 Scorecard metrics and their aggregated Scorecard score (computed by aggregating individual security practice scores) as predictors and Vul_Count, MTTR, and MTTU as target variables.
-Our initial dataset included 146K npm packages that have Scorecard, Vul_Count , and MTTU data.
-However, 85% lacked MTTR data because their dependencies either had no reported vulnerabilities or had vulnerabilities that remained unpatched.
-To address this, we extracted a 22K subset from the 146K dataset with MTTR data.
-Our findings show that higher aggregated Scorecard scores are associated with fewer Vul_Count and shorter MTTU, also supported by causal analysis.
-However, while the regression model suggests shorter MTTR, causal analysis indicates project characteristics likely influence MTTR direction.
-Segment analysis shows that larger, newer repositories with more contributors, dependencies, and downloads have shorter MTTR.
-Among individual security practices, Code Review, Maintained status, Pinned Dependencies, and Branch Protection show strong associations with security outcomes; the directionality of these associations varies across security outcomes.
-Additionally, MTTU follows trends similar to Vul_Count and MTTR, making it a potential proxy metric when vulnerability and MTTR data are unavailable.
+We also investigated project-level security outcome metrics: the number of open vulnerabilities (Vul_Count), mean time to remediate (MTTR) vulnerabilities in dependencies, and mean time to update (MTTU) dependencies.
+We conducted regression and causal analysis using 11 Scorecard metrics and the aggregated Scorecard score (computed by aggregating individual security practice scores) as predictors and Vul_Count, MTTR, and MTTU as target variables.
+Our findings reveal that aggregated adoption of security practices is associated with 5.2 fewer vulnerabilities, 216.8 days faster MTTR, and 52.3 days faster MTTU.
+Repository characteristics have an impact on security practice effectiveness: repositories with high security practices adoption, especially those that are mature, actively maintained, large in size, have many contributors, few dependencies, and high download volumes, tend to exhibit better outcomes compared to smaller or inactive repositories.
+Among individual security practices, Code Review, Contributors from diverse organizations, License, CI-Tests, and Pinned Dependencies show strong associations with security outcomes.
 1. Introduction
-The package release frequency has increased by 1,466% over the last decade, with smaller applications now incorporating an average of 180 components (sonatype2024, ).
-The growing reliance on third-party dependencies in modern applications expands the attack surface, exposing applications to CVE (Common Vulnerabilities and Exposures catalog) exploits or malicious code injection, facilitating software supply chain (SSC) attacks (williams2024research, ).
+The growing reliance on third-party dependencies in modern software applications expands the attack surface, exposing applications to CVE (Common Vulnerabilities and Exposures catalog) exploits or malicious code injection, facilitating software supply chain (SSC) attacks (williams2024research, ).
 For example, CVEs have grown by 463% in the last 10 years, and 704,102 malicious packages have been proactively identified since 2019 (sonatype2024, ).
-Practitioners are falling behind in managing vulnerabilities in dependencies due to growing dependency complexity and resource constraints, with mean time to remediate (MTTR) increasing from under 25 days in 2017 to over 400 days by 2024 (sonatype2024, ).
-SSC security standards have been proposed by industry and government agencies to help organizations adopt security practices that mitigate SSC security risks.
-Security practices are actions, procedures, techniques, or other measures that an organization can adopt to reduce the vulnerability of an information system (NISTSecurityControls, ).
-Examples of security practices include conducting code reviews before merging pull requests and using a static analyzer tool to detect vulnerabilities.
-In this paper, we used the term “security practice,” also synonymous with security control, safeguards, and countermeasures (NISTSecurityControls, ).
-Many stakeholders have recognized SSC security concerns; most prominently, the Executive Order (EO) 14028 (EO_2021, ) mandates the adoption and self-attestation of security practices to improve SSC security.
-Similarly, the European Union’s Cyber Resilience Act (CRA), approved in March 2024, mandates common cybersecurity standards for digital products, including incident reporting and automatic MTTR updates (ENISA_CRA_2024, ).
-In 2025, the US EO 14144 (EO2025Cybersecurity, ) strengthens the efforts by requiring- “Within 240 days of the date of this order,….the Director of NIST, shall evaluate common cybersecurity practices and security outcomes that are commonly used or recommended across industry sectors.” Security outcomes are the indicators of the performance or non-performance of security practices prescribed for an information system (EO2025Cybersecurity, ).
-While the SSC security standards (EO_2021, ; ENISA_CRA_2024, ; EO2025Cybersecurity, ; souppaya2022secure, ; Microsoft_framework, ; Williams2024PSSCRM, ) contain a long list of security practices that practitioners should adopt, practitioners are often overwhelmed by the number of practices recommended in these standards (zahan2023software, ).
-Given the limited budget, time, and resources (sammak2023developers, ; hamer2025closing, ), practitioners want to prioritize adopting security practices based on empirical evidence showing that adopting security practices mitigates security risk.
-The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices and security outcome metrics.
-Our research directly aligns with federal mandates EO 14144.
-However, instead of identifying commonly adopted security practices, we investigate the relationship between security practices and outcomes, providing empirical evidence to understand which practices are associated with improved security.
-We address the following research questions (RQs) and hypotheses:
-  • RQ1: Is higher adoption of security practices associated with measurable improvements in security outcomes?
+Practitioners are falling behind in managing vulnerabilities in dependencies due to growing dependency complexity and resource constraints, with mean time to remediate (MTTR) vulnerabilities increasing from under 25 days in 2017 to over 400 days by 2024 (sonatype2024, ).
+SSC security standards and frameworks (e.g.
+(EO_2021, ; ENISA_CRA_2024, ; EO2025Cybersecurity, ; souppaya2022secure, ; Microsoft_framework, ; Williams2024PSSCRM, )) have been proposed by industry and government agencies to help organizations adopt security practices that mitigate SSC security risks.
+However, historical experience demonstrates the risks of adopting security practices without empirical validation.
+For example, long password complexity requirements, such as mandating uppercase letters, numbers, and symbols, were long assumed to enhance security (adams1999users, ; morris1979password, ).
+However, empirical studies (dekoven2022measuring, ; dekoven2019measuring, ; lain2024content, ) later revealed that such policies often led users to create predictable passwords by encouraging reuse and minor modifications.
+Similar theoretical assumptions may persist in SSC security, where different practices are widely promoted without comprehensive empirical validation.
+Security success is often measured through adoption metrics—e.g., the number of teams that adopted code review or the number of vulnerabilities detected or fixed, rather than actual security outcomes.
+Therefore, little is known about the effectiveness of best practices in improving security.
+The absence of empirical evidence arises from multiple systemic challenges: widespread disagreement about best practices among experts (reeder2017152, ; redmiles2020comprehensive, ), poor evidence even where consensus exists (shostack2019empirical, ; thompson2017large, ), and dependence on intuition and convention rather than empirical evidence.
+The cybersecurity community’s reliance on perceived common sense has created an information-poor environment characterized by conflicting guidance and expert disagreement about fundamental priorities (reeder2017152, ; redmiles2020comprehensive, ; national2025cyber, ).
+Meanwhile, organizations have limited budgets, time, and resources (sammak2023developers, ; hamer2025closing, ), yet practitioners currently lack the empirical evidence needed to prioritize security practices that mitigate security risk.
+Therefore, developing empirical research to evaluate the security outcome of recommended practices is both a research priority and a fundamental prerequisite for evidence-based security decision-making (scott2025counting, ; national2025cyber, ).
+The goal of this study is to assist practitioners and policymakers in making informed decisions on which security practices to adopt by evaluating the relationship between software security practices adoption and security outcome metrics.
+We address the following research questions (RQs):
+  • RQ1: Is higher adoption of aggregated security practices associated with measurable improvements in security outcomes?
   • RQ2: Which security practices are most strongly associated with security outcomes?
-To that end, the study investigated the npm ecosystem to understand the relationship between publicly available data on security practice adoption and security outcomes.
-We selected the OpenSSF Scorecard (Scorecard, ) tool for our large-scale analysis because the tool utilizes metrics that automatically measure the adoption of security practices in the GitHub repositories.
-The Scorecard (Scorecard, ) uses 18 predefined metrics to detect the adoption of security practices.
+To that end, the study investigated the relationship between publicly available data on security practice adoption and security outcomes metrics.
+For large-scale analysis, we focus on security practice adoption that can be measured automatically.
+We selected the OpenSSF Scorecard tool (Scorecard, ) that automatically collects 18 metrics on the adoption of security practices in GitHub repositories.
+The Scorecard metrics are valuable for collecting ecosystem-scale data, as well as for evaluating the security posture of project dependencies.
+Examples of security practices that can be measured by Scorecard include maintainers conducting code reviews before merging pull requests, using a static analyzer tool, and others.
 For each metric, the Scorecard tool assigns a score of 0-10 based on pre-defined criteria.
-The tool also computes an aggregate security score as a weighted average of the 18 individual security metrics to understand the overall security of the repositories.
-For security outcomes, we selected the number of open vulnerabilities (Vul_Count), the mean time to remediate (MTTR), and the mean time to update (MTTU) metrics.
-To address RQ1, we developed a regression model to assess the relationship between a package’s aggregate security score and security outcome metrics.
-Additionally, we applied Propensity Score Matching (PSM), a statistical matching technique used in observational studies, to estimate causal effects and mitigate the potential confounding effects.
-For RQ2, we employed a regression model and feature importance-ranked techniques to identify which security practices have the strongest associations with security outcomes.
+The tool also computes an aggregate security score as a risk-weighted average of the individual security practice metrics, weighted by risk.
+For security outcomes, we selected project-based outcome metrics, where a project is defined as a software package that includes all components in its dependency graph.
+Specifically, the number of open vulnerabilities (Vul_Count), the mean time to remediate (MTTR) vulnerable and outdated dependencies, and the mean time to update (MTTU) outdated dependencies were used.
+To address RQ1, we investigate the aggregate adoption of security practices and their impact on security outcomes.
+We developed a regression model to assess the relationship using Scorecard aggregate security score and security outcome metrics.
+Additionally, we applied Propensity Score Matching (PSM), a statistical matching technique used in observational studies, to estimate causal effects and mitigate the potential confounding effects based on repository characteristics.
+For RQ2, we employed a regression model and feature importance-ranked techniques to identify which individual security practices have the strongest associations with security outcomes.
 To further validate these findings, we used PSM to estimate the effect of individual security practices on security outcomes, isolating their impact while accounting for differences in repository characteristics.
-Our research builds on Zahan et al.
-(zahan2023software, ), which investigated whether OpenSSF Scorecard metrics predict vulnerabilities in the PyPI and npm ecosystems.
-Their study found a counterintuitive positive association between security practices and vulnerabilities and attributed the result to a small dataset (2,422 packages).
-However, the authors collected vulnerabilities at the package level while obtaining security practices from the GitHub repositories, which often contain additional dependencies and languages, leading to incomplete vulnerability data.
-We improved data collection by 1) incorporating vulnerabilities from dependencies and repositories instead of package level; 2) constructing a large dataset (146k npm packages); 3) including control variables (project age, size, etc.) to address external factors that might have an influence; and 4) introducing MTTR and MTTU as novel security outcome metrics to provide a more comprehensive evaluation.
-The study will have the following contributions:
-  • MTTR and MTTU metrics to measure security outcomes.
+The study has the following contributions:
+  • Introduction of project-based security outcome metrics.
+  • Application of industry-standard metrics, MTTR and MTTU, to assess security outcomes.
   • A proposed model that can be used to explore the relationship between security practice and security outcome metrics.
-  • Empirical evidence on how security practices impact security outcome of npm packages.
-The paper is organized as follows: Section 2 discusses metrics used in our study, and Section 3 discusses our dataset construction steps.
-Section 4, Section 5 and Section 6 describes the data processing, methodology and result.
+  • Empirical evidence on how security practices impact the security outcome of npm packages.
+The paper is organized as follows: Section 2 contains the definition of security practice and security outcome; Section 3 discusses the design of our study; Section 3.1 discusses the metrics used in our study; and Section 3.2 discusses our dataset construction steps.
+Section 4, Section 5 and Section 6 describe the data processing, methodology, and results.
 We close with a discussion of our study (Section 7) and related work in Section 8.
-2. Metrics Utilized in the Study
+2. Definition
+Security practices are actions, procedures, and technical controls that an organization can adopt with the goal of preventing, detecting, and responding to threats against information systems.
+Security practices encompass both technical and organizational measures that are typically standardized and carried out by individuals or teams as part of an organization’s efforts to protect its information, resources, and operations (king2000best, ).
+Security outcomes are quantifiable indicators that reflect the effectiveness of security practices implemented within an information system (EO2025Cybersecurity, ).
+Outcome metrics assess whether implemented security practices are achieving their intended security risk management objectives, specifically reducing harm to information systems and guiding continuous security improvement.
+Security outcome metrics may include indicators such as reductions in attack frequency, security incidents, or financial losses; decreased vulnerability exposure (e.g., lower vulnerability counts), and faster response times, which shorten the window of exposure to threats.
+3. Experiment design
+In this section, we outline the experimental design and dataset used to investigate our research questions.
+To answer RQ1 and RQ2, we define Scorecard security practices metrics and aggregate security score as predictor variables and security outcome metrics as target variables.
+We also incorporate control variables to account for potential confounding project characteristics.
+We leveraged machine learning and causal analysis techniques to examine the relationship between practices and outcomes.
+Our analysis is based on data collected from npm packages and their corresponding GitHub repositories, using dependency graphs to define project scope.
+The dataset construction integrates multiple data sources and applies systematic inclusion and exclusion criteria to ensure reliability and reproducibility.
+Section 3.1 details the selection of metrics used in our study, while Section 3.2 describes how we collected the dataset.
+3.1. Metrics Utilized in the Study
 In this section, we discuss our metrics selection process, including security practices, security outcomes, and control variables.
-2.1. Security Practice Metrics
-In this study, we conducted a large-scale analysis to understand the adoption of security practices in OSS.
-We used the OpenSSF Scorecard tool (Scorecard, ), which evaluates security practices that can be measured automatically in GitHub.
-The tool detects security practice adoption, such as whether the package has implemented code review, repositories have enabled branch protection, or the repo is actively maintained.
+3.1.1. Security Practice Metrics
+To assess the adoption of security practices, the OpenSSF Scorecard tool (Scorecard, ) was employed.
+The tool automatically evaluates a set of security practices implemented in GitHub repositories, based on observable project metadata and activity.
 Table  1 contains the list of 18 security practices measured by Scorecard.
-Beyond industry adoption, the primary academic research on OpenSSF Scorecard has been conducted in study (zahan2023openssf, ; zahan2023software, ).
-The authors examine the adoption trends of security practices across ecosystems as measured by Scorecard metrics, analyze the distribution of Scorecard scores, and evaluate the efficacy of the Scorecard tool in assessing security practices.
-The author also verified whether the practices recommended by the Scorecard tool are also recommended in the standard security framework SSDF provided by NIST.
-The P-SSCRM SSC Framework (Williams2024PSSCRM, ) also included the OpenSSF Scorecard as one of the industry standards.
-2.2. Security Outcome Metrics
-Security outcomes are measured through quantitative characteristics, providing insights into the effectiveness of security practice adoption (scottmadden_outcome_metrics, ).
-In this study, we selected two temporal vulnerability metrics that capture both exposures (Vul_Count: number of open vulnerabilities) and responsiveness (MTTR: mean time to remediate vulnerable dependencies) and one temporal software metric (MTTU: mean time to update dependencies) as security outcome metrics.
-  • Vul_Count: A measure quantifying the number of publicly disclosed or reported open security vulnerabilities within a project and its dependencies.
+The tool also computes an aggregate security score to assess the overall security of a repository, computed as a risk-weighted average of individual security practice metrics.
+Scorecard was selected due to its widespread adoption and recognition in both academic and industry contexts.
+In academic research, Scorecard has been the subject of empirical evaluations, including SSC risk measurement (zahan2023software, ; mounesan2023exploring, ; butler2024links, ; salvievaluating, ; siddiqui2024elevating, ; kancharoendee2025categorizing, ; wu2025libvulnwatch, ; rydin2025evaluating, ), surveys on practitioners’ perspectives on Scorecard (elder2024applying, ; rydin2025evaluating, ), and qualitative studies (zahan2023openssf, ) examining its applicability and effectiveness.
+In industry, Scorecard has been adopted as part of standard security frameworks and requirements (Williams2024PSSCRM, ; openssf_osps_baseline_2025, ) and has also been used in reports by organizations, such as Sonatype and Veracode, to assess the state of open-source software ecosystems (sonatype_2022, ; eng2024quantifying, ).
+Adoption by major projects such as Kubernetes, Node.js, and Apache, along with integration into platforms like Ortelius (ortelius2024dashboard, ) and Deps.dev (OSI, ), also highlights wide adoption.
+3.1.2. Security Outcome Metrics
+A comprehensive understanding of security outcome metrics remains an open research challenge.
+Although defining and validating such metrics is inherently challenging, our work represents a step toward advancing that goal.
+In this study, we selected project-level outcome metrics because prior studies (zahan2023software, ) showed that only a small fraction of packages (0.1% of the npm ecosystem) have reported vulnerabilities at the individual package level.
+Limited availability of vulnerability data prevents models from learning meaningful relationships between security practices and security outcomes.
+Therefore, we leveraged three project-level outcome metrics, where a project is defined as an individual package and its dependencies.
+The security risk of a project decreases when a vulnerable component is replaced with a fixed one, providing a measurable indication of improved security outcomes, reflected in fewer open vulnerabilities and reduced exposure windows due to faster response times.
+  • Vul_Count: A measure quantifying the number of publicly disclosed or reported open security vulnerabilities within a package and its full dependency tree, including both direct and transitive dependencies.
   • MTTR: A measure of the average aggregated time a package uses outdated and vulnerable direct dependencies in its lifetime (rahman2024characterizing, ).
@@
 The number of vulnerabilities is widely used in academic research, industry, and policy as outcome metrics (EO2025Cybersecurity, ; CISASecureSoftwareAttestation, ; morrison2014mapping, ; sec_outcome, ; morrison2015security, ; zahan2023software, ).
-MTTR is most commonly used to measure the package’s security (MTTR_1, ; MTTR_2, ), measuring the code dependencies security (MTTR_3, ) in the industry.
-The OpenSSF Secure Supply Chain Consumption Framework (S2C2F) framework (Microsoft_framework, ) includes MTTR as a key maturity indicator, while Sonatype (sonatype-2019, ) has used MTTR to classify secure OSS projects.
-However, measuring MTTR poses a challenge since many package dependencies either have no reported vulnerabilities or remain unpatched, limiting the dataset size.
-To mitigate this limitation, we also consider MTTU to be a proxy metric for MTTR.
-Rahman et al.
-(rahman2024characterizing, ) found a moderate correlation between MTTR and MTTU, suggesting that MTTU can used as a proxy when MTTR data is unavailable.
-We consider MTTU as an alternative indicator of MTTR: 1) MTTU data was derivable for all components; and 2) We hypothesize developers staying up to date on dependencies will also stay up-to-date on security updates because security updates are a subset of regular updates.
-Hence, packages will exhibit similar MTTR and MTTU values.
-The hypothesis, data availability, and the observed correlation between MTTR and MTTU (rahman2024characterizing, ) motivated us to consider MTTU as an additional security outcome metric.
-2.3. Control Variable Metrics
-Control variables are properties that researchers hold constant for all observations in an experiment to understand the true relationships between the predictor and target variables (FrostControlVariables, ).
-In this research, we evaluated the impact of security practices on security outcomes while including other factors that may also influence the outcome.
-We included five control variables to measure project characteristics.
-GitHub contributions count controls for developer activity, while package downloads account for package popularity.
-Repository age (years) controls the project’s development stage, and size (KB) controls the complexity of maintaining a larger package.
-Direct dependency count controls a project’s reliance on and exposure to external dependencies.
+While MTTR and MTTU are widely used industry metrics, our work is novel in applying them as outcome measures for security responsiveness.
+Faster updates reduce exposure windows, limiting attackers’ opportunity to exploit disclosed weaknesses.
+We draw inspiration from the DevOps Research and Assessment (DORA) framework (dora_metrics, ).
+DORA is one of the largest research studies in software engineering research, analyzing survey responses from over 31,000 professionals across six years (researched in the years 2014 through 2019) (forsgren2019accelerate, ).
+DORA research identifies which software engineering practices, including security practices, improve software delivery performance.
+One of DORA’s outcome metrics is a responsiveness metric that measures how long it takes a project to respond and deploy the corresponding updates.
+Research conducted in 2022 by the Google DORA team (google2022accelerate, ) specifically investigated SSC practices derived from the Secure Software Development Framework (SSDF) (souppaya2022secure, ) and Supply-chain Levels for Software Artifacts (SLSA) (SLSA, ), and examined their impact on outcomes.
+The research has shown that SSC practices are positively associated with faster response and deployment times, and practitioners also anticipate a lower chance of security vulnerabilities.
+Motivated by DORA metrics, we hypothesize that teams that adopt good security practices are also more likely to be more responsive, as measured by MTTR and MTTU.
+We used MTTU because ecosystem-wide MTTR data is often unavailable (only 15% of our dataset; see section 3.2.4).
+While MTTR focuses on the response to security updates, MTTU measures responsiveness to any upstream updates and is available for all packages.
+Prior work (rahman2024characterizing, ) also found a moderate correlation between MTTR and MTTU, suggesting that MTTU can serve as a proxy when MTTR is unavailable.
+3.1.3. Control Variables
+Control variables are properties that researchers hold constant in an experiment to understand the true relationships between the predictor and target variables (FrostControlVariables, ).
+We included six variables to measure repository characteristics based on two criteria: (1) alignment with prior use of control variables on software quality and practices work (miller2025understanding, ; zhou2019fork, ; ray2014large, ; ortu2019empirical, ; trockman2018adding, ; li2024can, ; javan2023dependency, ), and (2) maintaining model interpretability with a minimal yet effective set of metrics to reduce data collection and cost.
+We controlled for six repository characteristic metrics grouped by their hypothesized influence.
+Contributor count (Contributors_CT) and commit staleness reflect team capacity and development activity, which may affect the implementation and maintenance of security practices.
+Download count and direct dependencies capture project adoption and external exposure, potentially influencing both scrutiny and security risk.
+Repository age and size represent project maturity and complexity, which may impact maintenance effort and security performance.
+We excluded highly correlated variables such as forks, stars, and dependents due to redundancy and because they did not improve model performance.
 Table 1. Scorecard Security Practices Metrics
-Rows with bolded security practice names indicate excluded metrics, either due to high computational expense (HCE) or because more than 90% of repositories had a score of -1.
-A score of -1 indicates that Scorecard could not get conclusive evidence of implementing practices, or perhaps an internal error occurred due to a runtime error.
-A score between 0-10 represents the extent of adoption of the security practice.
-Security Practices Description Risk Label -1 score 0-10 score
-Dangerous-Workflow Does the project avoid dangerous coding patterns in GitHub Action workflows? Critical 73.2% 26.8%
-Branch-Protection Does the project use Branch Protection? High ¡1% (7) 99.9%
-Code-Review Does the project practice code review before code is merged? High ¡1.4% (2071) 98.6%
-Binary-Artifacts Is the project free of checked-in binaries? High ¡1% (3) 99.9%
-Dependency-Update-Tool Does the project use tools to help update its dependencies? High HCE HCE
-Maintained Is the project at least 90 days old and maintained? High 0 100%
-Signed-Releases Does the project cryptographically sign releases? High 97.7% 2.3%
-Token-Permissions Does the project declare GitHub workflow tokens as read only? High 74.4% 26.8%
-Vulnerabilities Does the project have unfixed vulnerabilities (including depedencies)? High 0 100%
-Pinned-Dependencies Does the project declare and pin dependencies? Medium 72.2% 27.8%
-Packaging Does the project build and publish official packages from CI/CD, e.g. GitHub Publishing ? Medium 97.4% 2.5%
-Fuzzing Does the project use fuzzing tools, e.g. OSS-Fuzz, QuickCheck or fast-check? Medium ¡1% (15) 99.9%
-Static Application Security Testing (SAST) Does the project use static code analysis tools, e.g. CodeQL, LGTM (deprecated), SonarCloud? Medium ¡1% (203) 99.9%
-Security-Policy Does the project contain a security policy? Medium ¡1% (3) 99.9%
-CI-Tests Does the project run tests in CI, e.g. GitHub Actions, Prow? Low HCE HCE
-Contributors Does the project have contributors from at least two different organizations? Low HCE HCE
-License Does the project declare a license? Low ¡1% (2) 99.9%
-CII-Best-Practices Has the project earned an OpenSSF (formerly CII) Best Practices Badge at the passing, silver, or gold level? Low ¡1% (12) 99.9%
-3. Dataset Construction
-The primary source of our dataset are the npm packages and their dependency graph and GitHub repositories.
-In Section 3.1, we discuss different sources used to construct our dataset.
-The package inclusion and exclusion criteria are discussed in Section 3.2.
-Section  3.3 discusses our security practice metrics data collection process using the OpenSSF Scorecard tool.
-Section 3.4 and Section 3.5 cover our security outcome metrics and control variables.
-3.1. Data Source
+-1: no conclusive evidence or runtime error. 0–10: extent of security practice adoption. Values in parentheses indicate raw counts.
+Security Practices Description -1 % 0% 1–10%
+Binary-Artifacts Is the project free of checked-in binaries? 0.0 0.2 99.8
+Branch-Protection Does the project use Branch Protection? 0.0 (3) 85.9 14.1
+Code-Review Does the project practice code review before code is merged? 0.4 75.4 24.2
+License Does the project declare a license? 0.0 30.9 69.1
+Maintained Is the project at least 90 days old and maintained? 0.0 86.7 13.3
+Dependency-Update-Tool Does the project use tools to help update its dependencies? 0.0 (65) 81.3 18.7
+Contributors Does the project have contributors from different organizations? 0.0 36.7 63.3
+SAST Does the project use static code analysis tools, e.g. CodeQL? 0.0 (23) 96.9 3.1
+Security-Policy Does the project contain a security policy? 0.0 (2) 93.2 6.8
+CI-Tests Does the project run tests in CI, e.g. GitHub Actions, Prow? 50.3 41.3 8.4
+Pinned-Dependencies Does the project declare and pin dependencies? 72.2 18.0 9.8
+Token-Permissions Does the project declare GitHub workflow tokens as read-only? 73.2 25.6 1.2
+Dangerous-Workflow Does the project avoid dangerous patterns in GitHub workflows? 73.2 0.3 26.6
+Fuzzing Does the project use fuzzing tools, e.g. OSS-Fuzz, QuickCheck? 0.0 (2) 99.8 0.2
+CII-Best-Practices Has the project earned an OpenSSF Best Practices Badge? 0.0 99.9 0.1
+Signed-Releases Does the project cryptographically sign releases? 97.7 2.3 0.0
+Packaging Does the project build and publish official packages from CI/CD? 97.5 0.0 2.5
+Vulnerabilities Does the project have unfixed vulnerabilities (including dependencies)? 0.0 32.2 67.8
+3.2. Dataset Source and Construction
+Section 3.2.1 discusses different sources used to construct our dataset.
+Section 3.2.2, Section 3.2.3, Section 3.2.4 and Section 3.2.5 cover our dataset collection process.
+Our dataset is available in Figshare [^1] https://figshare.com/s/e74f94dfe1be6a90c4d1?file=54861686.
+3.2.1. Data Sources
 To collect data, we leveraged multiple sources.
@@
 Within Scorecard’s dataset of selected popular packages, 64% are npm packages.
-Additionally, our security outcome metrics require analyzing project dependencies, specifically counting vulnerabilities in dependencies and measuring dependency update patterns for MTTR and MTTU.
+Additionally, our security outcome metrics require analyzing project dependencies, specifically counting vulnerabilities in dependencies and measuring outdated and vulnerable dependency update patterns for MTTR and MTTU.
 To support our analysis, we collected the dependency graph of npm packages and mapped packages to their corresponding GitHub repositories using deps.dev (OSI, ).
 Deps.dev, developed and hosted by Google, provides comprehensive insights into OSS, including source code locations, dependency graphs, package metadata, licenses, releases, and vulnerabilities.
-Another platform used in our study is OSV (Open Source Vulnerabilities) (osv_adv, ).
-The OSV platform is an aggregator of vulnerabilities from different security advisories.
+Another platform used in our study is Open Source Vulnerabilities (OSV) (osv_adv, ).
+The OSV platform aggregates vulnerabilities from various security advisories.
 Scorecard and Deps.dev depend on OSV to collect vulnerability information for projects and their dependencies.
-We also used the Libraries.io platform (LibrariesIO, ), which collects publicly available OSS information scraped from the internet, including GitHub.
-3.2. Package Inclusion and Exclusion Criteria
+We also used the Libraries.io platform (LibrariesIO, ), which collects publicly available OSS metadata scraped from the internet, including GitHub.
+3.2.2. Package Inclusion and Exclusion Criteria
 To construct our dataset, we systematically filtered, cleaned, and processed npm packages based on dependency relationships, repository availability, and data availability.
@@
   • Initial Dataset: We began with a total of 2,119,044 unique npm packages extracted from the Deps.dev BigQuery dataset (depsdevBigQuery, ).
-This dataset contains the mapping between npm packages to GitHub repositories and captures all dependency edges within the npm.
-  • Packages with GitHub Repositories: Of these, 1,241,732 packages were mapped to GitHub repositories.
-Since we must map each package to a Scorecard score via a GitHub repo, we exclude 877,312 packages that lack repository information, either due to hosting outside GitHub or missing repository URLs.
+This dataset contains the mapping between npm packages and GitHub repositories and captures all dependency edges within npm.
+  • Packages with GitHub Repositories: Of these, 877,312 packages were excluded because they lacked valid GitHub repository information, either because they were hosted outside GitHub or had missing repository URLs.
+Mapping packages to GitHub repositories was necessary because Scorecard runs directly on GitHub repositories.
+After this filtering step, 1,241,732 packages remained that could be successfully mapped to GitHub.
   • Packages with Dependencies and Dependents: Further filtering identified 1,159,891 packages with at least one declared dependency, ensuring relevance to our security outcome metrics.
-Additionally, we are primarily interested in packages that practitioners may decide to use; hence, we refined the dataset further to include only packages that had at least one dependent, reducing the dataset to 264,413 packages.
+Additionally, we are primarily interested in packages that practitioners may use; hence, we refined the dataset further to include only packages with at least one dependent, reducing the dataset to 264,413 packages.
   • Ensuring Unique Repositories: Scorecard generates metric scores based on security practice adoption at the repository level rather than for individual packages.
-While we collected vulnerability data at the repository level (Section 3.4), MTTR and MTTU can only be measured at the package level.
+While we collected vulnerability data at the repository level (Section 3.2.4), MTTR and MTTU can only be measured at the package level.
 Consequently, if a repository contains multiple packages, all packages share the same Scorecard score and vulnerability data but may have different MTTR and MTTU values.
@@
 After applying this filtering step, we retained 172,999 unique repositories.
-3.3. OpenSSF Scorecard Data
-Once we identified our set of 172,999 unique npm GitHub repositories, we collected OpenSSF Scorecard scores for each repository associated with a package.
+3.2.3. OpenSSF Scorecard Data
+Once we identified the set of 172,999 unique npm GitHub repositories, we collected OpenSSF Scorecard scores for each repository associated with a package.
 Our primary data source was the OpenSSF Scorecard dataset in Google BigQuery (ScorecardBigQuery, ).
 However, Scorecard only evaluates a subset of repositories, specifically popular packages and those explicitly requested by practitioners for inclusion.
-From Google BigQuery, we extracted Scorecard data for 761,602 repositories.
-However, not all of these repositories met our inclusion criteria mentioned in Section 3.2.
-For example, some lacked dependencies or dependents, while others had multiple packages mapped to single repositories, which was outside the scope of our study.
 Hence, many repositories in our dataset were not included in the original Scorecard dataset.
 To address these gaps, we ran Scorecard locally for any repositories missing from the BigQuery dataset, using the same configuration applied to populate BigQuery.
-After merging data from Google BigQuery and our local runs, we obtained Scorecard scores for 146,951 out of 172,999 repositories.
+After merging data from Google BigQuery and our local runs, we obtained Scorecard scores for 145,817 out of 172,999 repositories.
 The remaining repositories did not generate a score using Scorecard.
-Upon manual verification of 100 repositories, we found that Scorecard could not produce scores primarily because the repositories were either private or had been removed from GitHub.
-Metrics inclusion and exclusion In this study, we studied 13 out of 18 Scorecard metrics, excluding CI-Tests, Contributors, Dependency-Update-Tool, Signed Releases, and Packaging, due to computational constraints or a lack of sufficient data.
-The CI-Tests, Contributors, and Dependency-Update-Tool metrics are not included in the Google BigQuery dataset due to the computational resources required to run these checks, which would hinder periodic updates.
-The exclusion was confirmed by the OpenSSF team, who specified the configuration in the project’s source code  [^1] https://github.com/ossf/scorecard/blob/76a04bfe40c25746b2633127270a9219a98a37ad/cron/config/config.yaml#L48.
-Scorecard generates a -1 score when no conclusive evidence of a practice’s adoption exists, either due to insufficient repository data or internal runtime errors.
-Signed Releases and Packaging were excluded, as over 90% of the repositories received a -1 score (Table 1).
+Upon manual verification of a random sample of 100 repositories and their Scorecard output, we found that Scorecard failed to produce scores due to HTTP 404 errors.
+The primary reason for the lack of scores was that the repositories were no longer accessible via public GitHub URLs.
+The repositories were either private or had been removed from GitHub.
+Metrics exclusion: We began by collecting all 18 metrics provided by Scorecard.
+However, due to limited adoption and Scorecard measuring limitations, we excluded Fuzzing, CII-Best-Practices, Signed-Releases, and Packaging from our analysis (Table 1).
+Fuzzing and CII Best Practices were excluded because  99.8% of repositories scored 0, indicating minimal adoption of Scorecard recommendations across ecosystems.
+Signed Releases and Packaging were excluded, as almost 98% of the repositories received a -1 score due to not having any signed release or a package release workflow in GitHub.
 Prior research (zahan2023openssf, ) found that practitioners primarily used package registries for releases rather than GitHub, and no improvements in Scorecard’s methodology addressing this limitation were observed.
+We also removed the vulnerability metric as a practice because we use Vul_Count as a dependent variable.
 Thus, we removed five metrics from the analysis.
-Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions metrics have less than 75% repositories with a -1 score (Table 1).
-Prior research by Zahan et al.
-(zahan2023openssf, ) highlighted limitations in Scorecard’s tool, particularly for these metrics that are used to assess security configuration in GitHub workflows and dependency management practice (Table 1).
-Previously, Scorecard assigned a maximum score of 10 for these metrics, even when no GitHub workflows or dependency metadata were present, which could mislead practitioners about positive security practice adoption.
-Our manual analysis of 50 repositories per metric confirms that Scorecard has improved the detection of these practices.
-Scorecard assigns a -1 score when repositories do not have any GitHub workflows, token information, or dependency metadata.
-A score of 0 now signifies that the practice exists with security misconfiguration, such as excessive GitHub token permissions, dangerous workflow patterns, or all dependencies are unpinned.
-We included these three metrics in our study because a -1 score represents non-adoption, while a score of 0 indicates security misconfigurations.
-3.4. Security Outcome Metrics Data
-We collected our security outcome metrics for 146,951 npm packages with Scorecard data. Here, we describe our outcome metrics data collection process.
-Vul_Count metric: The Scorecard vulnerability metrics measure whether the project has open, unfixed vulnerabilities in its codebase or its dependencies using the OSV (osv_adv, ) service.
-Initially, we used Deps.dev BigQuery (depsdevBigQuery, ) and the Scorecard tool to collect the total vulnerability count.
-Deps.dev (OSI, ) includes Scorecard metrics for OSS.
-In deps.dev, we were able to map vulnerability data for 76,651 packages, and 70,300 remaining data were collected using Scorecard.
-The scorecard normalizes the vulnerability score, where a score of 10 indicates no vulnerabilities, 9 corresponds to one vulnerability, and 0 represents more than nine vulnerabilities.
+3.2.4. Security Outcome Metrics Data
+We collected our security outcome metrics for 145,817 npm packages with Scorecard data. Here, we describe our outcome metrics data collection process.
+Vul_Count metric: The Scorecard Vulnerability metric measures whether the project has open, unfixed vulnerabilities in its codebase or its dependencies using the OSV (osv_adv, ) service.
+Initially, we used the Scorecard tool to collect the total vulnerability count.
+The Scorecard tool normalizes the vulnerability score, where a score of 10 indicates no vulnerabilities, 9 corresponds to one vulnerability, and 0 represents more than nine vulnerabilities.
 To obtain the exact vulnerability count, we first denormalized the score.
-For cases where the vulnerability score was 0 (indicating more than nine vulnerabilities), we ran the vulnerability metric separately using Scorecard to retrieve the raw, unnormalized data.
+For cases where the vulnerability score was 0 (indicating more than nine vulnerabilities), we ran the vulnerability metric separately using Scorecard to retrieve the raw, unnormalized data and used the total number of vulnerabilities in the model.
 MTTR and MTTU metric: In this study, we adopt the algorithms proposed by Rahman et al.
 (rahman2024characterizing, ) for computing MTTR and MTTU.
-Following their methodology, we contacted Google deps.dev team and considered deps.dev as our primary data source to collect comprehensive package versions and dependency versions and their release information for our dataset.
+Following their methodology, we contacted the Googlee deps.dev team and considered deps.dev as our primary data source to collect comprehensive package versions, dependency versions and their release information for our dataset.
 Additionally, we retrieved CVE data for selected packages and their direct dependencies from OSV (osv_adv, ).
 Note that while MTTU captures all dependency updates, MTTR specifically accounts for updates addressing vulnerable dependencies.
-While we successfully collected Vul_Count and MTTU data for 146,951 packages, MTTR data was only available for 22,083 packages (15% of total dataset).
-The limitation arises because MTTR penalizes packages that did not update their dependencies despite a known vulnerability being fixed in an upstream version.
-3.5. Control Variable Metrics Data
+MTTR needs a vulnerability in direct dependency, then the availability of a new fix version, and the adoption time to adopt the new fix version by the dependent.
+While we successfully collected Vul_Count and MTTU data for 145,817 packages, MTTR data was only available for 22,412 packages (15% of total dataset).
+3.2.5. Control Variable Metrics Data
 We collected control variable data from different sources.
-For example, GitHub contribution count and package size are collected from Libraries.io (LibrariesIO, ).
+GitHub contribution count and package size are collected from Libraries.io (LibrariesIO, ).
 Then we got the number of downloads of a package in the past 12 months, which we collected from the public npm API [^2] https://api.npmjs.org/downloads/point/{period}[/{package}].
-We calculated the repository age (in years) using the GitHub REST API (GitHubAPI2024, ) by retrieving the repository creation date and computing the difference between the current date and the creation date.
+We calculated the repository age (in years) and commit staleness dates using the GitHub REST API (GitHubAPI2024, ).
+The number of dependency information is collected from Deps.dev (depsdevBigQuery, ).
 4. Data Preprocess
-Before constructing machine learning (ML) modeling to answer RQ1, we performed data preprocessing steps to handle missing values, ensure data integrity, and assess statistical properties.
-Section 4.1 covers data imputation techniques to manage missing values.
+Before constructing the machine learning (ML) modeling to answer RQs, we performed data preprocessing steps to handle missing values, ensure data integrity, and assess statistical properties.
+Section 4.1 covers feature decomposition and imputation techniques to manage missing values.
 Section 4.2 discusses correlation and multicollinearity in our dataset.
-Figure 1 highlights the overview of our research workflow.
-Figure 1. Research Workflow
-4.1. Data Imputation
-Data imputation is the process of replacing missing values to maintain data integrity, prevent biases in ML models, and ensure statistical validity (schafer2002missing, ).
-Imputation techniques help mitigate data loss, improve model accuracy, and prevent misleading conclusions caused by missing data.
-Depending on the nature of the missing data, multiple imputation techniques exist, such as mean/median imputation, categorical encoding, and predictive modeling-based imputations.
-In our dataset, Branch-Protection, Code-Review, Binary-Artifacts, Fuzzing, SAST, Security-Policy, License, and CII-Best-Practices each had less than 1% of packages with a -1 score.
-Table 1 shows the exact counts of packages with -1 for these metrics.
-On the contrary, Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions had greater than 72% of repositories with a -1 score (Table 1).
-Our manual analysis found that Scorecard assigned -1 in two cases: (1) when there was no conclusive evidence of security practice adoption, and (2) due to internal runtime errors (section 3.3).
-Since not all metrics within a repository had a -1 score, removing entire repositories would result in unnecessary data loss.
-To handle missing data effectively, we differentiated between cases where -1 indicated non-adoption (e.g., missing GitHub workflows, token permissions, or dependency management practices) and cases where -1 resulted from internal runtime errors.
-For metrics where -1 was due to internal errors, we applied median imputation to maintain data completeness while minimizing distortion.
-For Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions, where -1 indicated non-adoption rather than missing data, we applied categorical encoding imputation to preserve the distinction.
-Note that a score of 0 reflects detected security misconfigurations.
+4.1. Feature Decomposition and Imputation
+Our manual analysis found that Scorecard assigned -1 in two cases: (1) when there was no conclusive evidence of security practice adoption; and (2) due to internal runtime errors.
+Since not all metrics within a repository had a -1 score (Table 1), excluding an entire repository due to a -1 in any one of the 18 metrics would result in unnecessary data loss.
+To handle missing data effectively, we implemented a missing data pattern decomposition strategy based on Scorecard score reasoning.
+For high-missing features CI-Tests, Pinned-Dependencies, Token-Permissions, and Dangerous-Workflow (each with $>$50% -1 scores), a score of -1 indicates that the repository lacks the relevant configuration.
+Here, -1 for CI-Tests indicates no pull requests to confirm CI workflows were present; for Pinned-Dependencies, it suggests the absence of declared dependencies in GitHub; for Token-Permissions, it reflects that no GitHub token settings were configured; and for Dangerous-Workflow, -1 means no GitHub Actions workflows were found to test workflow patterns.
+To preserve the distinction, we decomposed each feature into two components: a binary indicator of relevant configuration presence (1 for values 0-10, 0 for -1) and scores that capture adoption effectiveness (keeping original discrete values 1-10, with both -1 and 0 scores mapped to 0).
+For low-missing features ($<$1% -1 values), where -1 was due to internal errors of Scorecard, we applied median imputation (schafer2002missing, ) to maintain data completeness while minimizing distortion.
 Figure 1. Spearman Correlation between Predictors
 4.2. Correlation Analysis
-To evaluate dataset normality, we applied the Henze-Zirkler multivariate normality test(henze1990class, ) separately to the RQ1 and RQ2 datasets  (Figure 1).
-The RQ1 dataset includes aggregated Scorecard score, Control Variables, Vul_Count, MTTR, and MTTU, while RQ2 leverage 12 Scorecard Metrics instead of aggregated Scorecard score.
-Both datasets failed the normality test (p ¡ 0.05), indicating deviations from a normal distribution.
-To further assess skewness, we flagged features with —skewness— ¿ 1 for potential transformation (hatem2022normality, ).
-Based on this, we applied log(x + 1) transformation to highly skewed variables (e.g., GitHub contributors, repository size, DependencyCount, downloads, MTTU, and Vul_Count) to mitigate skewness.
+To evaluate dataset normality, we applied the Henze-Zirkler multivariate normality test (henze1990class, ) separately to the RQ1 and RQ2 datasets.
+The RQ1 dataset includes aggregate scorecard scores, the control Variables, Vul_Count, MTTR, and MTTU.
+The RQ2 dataset contains 13 Scorecard practice metrics, control variables, and Vul_Count, MTTR, and MTTU.
+Both datasets failed the normality test (p $<$ 0.05), indicating deviations from a normal distribution.
+To further assess skewness, we flagged features with $|skewness|>$ 1 for potential transformation (hatem2022normality, ).
+Based on skewness, we applied a log(x + 1) transformation to highly skewed control variables (GitHub contributors, repository size, DependencyCount, downloads) to mitigate skewness.
 Post-transformation, Q-Q plots, histograms, and statistical tests confirmed improved distribution symmetry.
+We did not apply transformation to the practice or outcome metrics to preserve their original scale and interoperability.
 Since residual normality is important for regression analysis, we conducted D’Agostino-Pearson tests (keskin2006comparison, ) on regression residuals.
-Results showed that residuals remained non-normally distributed (p ¡ 0.05).
+Results showed that residuals remained non-normally distributed (p $<$ 0.05).
 Given the non-normal distribution of our dataset, we proceeded with the Spearman correlation (zar2005spearman, ), a nonparametric method suitable for capturing monotonic relationships in non-normal data.
-We conducted correlation analysis for predictor variables (security practice and control variables) and target variables (security outcome metrics).
-Figure  2 presents the Spearman correlation heatmap for predictor variables, revealing weak to moderate correlations.
-The highest correlation is observed between contributors and code review (0.61), suggesting that repositories with more contributors are more likely to implement code review practices.
+We conducted a correlation analysis for predictor variables (security practice and control variables) and target variables (security outcome metrics).
+Figure  1 presents the Spearman correlation heatmap for predictor variables.
+The highest correlations ( 1) were observed among Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions.
+To avoid multicollinearity and redundancy in the model (guyon2003introduction, ), we removed Dangerous-Workflow and Token-Permissions, retaining only Pinned-Dependencies as a representative feature.
 The correlation between target variables reveals a strong positive correlation (0.7) between MTTR and MTTU, aligning with our expectation of potential overlap (rahman2024characterizing, ).
-In contrast, the Vul_Count metric exhibits a weak negative correlation with MTTR (-0.32) and MTTU (-0.04).
-Hence, we confirmed that MTTR and MTTU cannot be used as target variables in the same model due to their strong correlation since including both would introduce redundancy, bias, and multicollinearity.
-Additionally, we computed the Variance Inflation Factor (VIF) (o2007caution, ) for all predictor variables to confirm multicollinearity.
-For VIF computation, we found only “downloads” metrics had a high VIF score (15).
-According to study (o2007caution, ), VIF thresholds should be interpreted in context rather than strictly applied as exclusion criteria.
-Since the ”download” metric serves as a control variable in our study and does not exhibit strong correlations (¿0.8) with any other metrics, the high VIF alone is not sufficient grounds for removal (o2007caution, ).
+In contrast, the Vul_Count metric exhibits a weak negative correlation with MTTR (-0.3) and MTTU (-0.4).
+Hence, we confirmed that MTTR and MTTU cannot be used as target variables in the same model due to their strong correlation, since including both would introduce redundancy, bias, and multicollinearity.
+Additionally, we also computed the Variance Inflation Factor (VIF) (o2007caution, ) for all predictor variables to confirm multicollinearity.
+For the VIF computation, we did not find any feature that exhibited a VIF score greater than 5, indicating no multicollinearity.
 5. RQ1: Overall security posture and its security outcomes
-In this section, we discuss our RQ1: Is higher adoption of security practices associated with measurable improvements in security outcomes?
-Section 5.1 discusses our methodology to answer our RQ1, and Section 5.2 discusses our result.
-Table 2. Summary Statistics of Security Outcomes
-Statistic Vul_Count MTTR MTTU
-Count 146,951 22,543 146,951
-Mean 16.4 1,084.7 634.3
-Std Dev 31.3 897.3 877.4
-Min 0 1 0
-Median 1.0 940.0 222.2
-Max 637.0 4,709.0 4,869.6
+In this section, we discuss our RQ1: Is higher adoption of aggregated security practices associated with measurable improvements in security outcomes?
+Section 5.1 discusses the methodology of RQ1, and Section 5.2 discusses the results.
+Figure 2. Security Outcome Trends and npm Repository Distribution Across Aggregated Scorecard Score Bins
 5.1. RQ1: Methodology
-For our RQ1, we evaluate the relationship between npm package repositories’ aggregated Scorecard score and control variables as predictors and security outcome metrics as the target variable (Figure 1).
-Table 2 presents the summary statistics of security outcomes.
+For RQ1, we evaluate the relationship between the aggregated Scorecard score of npm package repositories and control variables as predictors, and security outcome metrics as the target variable.
 To visualize the relationship between aggregated Scorecard scores and security outcomes, we first binned repositories by aggregated Scorecard scores and analyzed mean security outcomes within each bin (Figure 2).
-Next, we applied regression analysis to quantify the association of the aggregated Scorecard score on outcomes.
-Finally, we conducted a causal analysis to assess whether the observed relationships reflect a causal effect.
+Next, we applied regression analysis to quantify the association of the aggregated Scorecard score with outcomes.
+However, regression analysis does not establish causality.
+We implemented a Propensity Score Matching (PSM) framework with caliper-based matching, covariate balance diagnostics, and segment-level analysis to estimate the causal effect of security practice adoption on outcome metrics, controlling for project characteristics.
+5.1.1. Data Expolation
+To visualize initial relationships, we binned repositories by aggregated Scorecard scores into five equal-width bins and calculated mean security outcomes within each bin.
+The binning approach provides an intuitive view of how security outcomes trend across different score ranges before applying statistical modeling.
 5.1.2. Regression Model:
-We selected Generalized Linear Models (GLMs) (nelder1972generalized, ), given the non-normal distribution of our dataset and our target variables are count data.
-GLMs are specifically designed for count data and provide interpretable coefficients.
+At this stage, we focus on interpreting how security outcomes change per unit change in the aggregated Scorecard score (e.g., a coefficient of -0.1 implies a 10% decrease in the outcome for every one-unit increase in aggregated Scorecard score).
+We selected Generalized Linear Models (GLMs) (nelder1972generalized, ), given the non-normal distribution of our dataset, and our target variables are count data.
+GLMs are explicitly designed for count data and provide interpretable coefficients.
 Coefficients quantify the expected change in the target variable for a unit increase in a predictor.
-To identify the best-fitting GLM, we systematically evaluated the dataset characteristics using a data-driven approach.
-Our selection process was guided by two factors: the Overdispersion Test (variance greater than the mean) and the zero-inflation test(a large proportion of zero values in target variables).
-We first started with the Poisson Regression (hausman1984econometric, ) as our baseline model and found that the dataset contains overdispersion for Vul_Count , MTTR, and MTTU, and zero inflation exists for Vul_Count only.
-We selected Negative Binomial (NB) Regression (hilbe2011negative, ) for MTTR and MTTU to address overdispersion and Zero-Inflated Negative Binomial (ZINB) (ridout2001score, ) to handle both overdispersion and zero inflation for Vul_Count.
-Since GLMs inherently apply a log link function instead of log transformation, we used the raw target variables to maintain the interpretability of the coefficients.
-Figure 3. Security Outcome Trends and npm Repository Distribution Across Aggregated Scorecard Score Bins
-5.1.2. Causal Analysis:
-While regression models can identify associations between “aggregated Scorecard score” and security outcomes, they do not establish causality due to the potential influence of confounding factors.
-For example, the predictors aggregated Scorecard score and control are correlated, which could bias the estimation of the true impact of security practices on security outcomes.
-To address this, we used Propensity Score Matching (PSM) (austin2009balance, ), a statistical technique that simulates a randomized experiment by pairing repositories with similar characteristics.
-PSM matched the treated units (repositories with aggregated Scorecard scores above the median) with control units (below the median) based on covariates.
-Covariates are control variables that could influence both aggregated Scorecard scores and security outcomes.
-Using logistic regression, we estimated each repository’s likelihood of having a high aggregated Scorecard score.
-We then applied nearest-neighbor matching with a caliper adjustment(0.1) to pair treated repositories with control repositories based on similar covariates.
-The similarity of matching was determined using propensity scores, which represent the likelihood that a repository belongs to the treated group based on the repository characteristics.
-To validate the matching process, we used standardized mean differences (SMD) to assess balance across covariates.
-Finally, we calculated the Average Treatment Effect (ATE) using an inverse probability-weighted (IPW) estimate (ATE weighted) that adjusts for any residual differences.
-Computing the Average Treatment Effect (ATE), we estimated how much a package’s security outcome would change if it had a higher aggregated Scorecard score.
+To identify the best-fitting GLM, we systematically evaluated dataset characteristics using standard diagnostic tests: overdispersion (variance $<$ mean) and zero-inflation.
+We began with Poisson Regression (hausman1984econometric, ) as our baseline model and found significant overdispersion across all outcomes, with zero-inflation present only in Vul_Count.
+We applied Negative Binomial (NB) Regression (hilbe2011negative, ) to address overdispersion and attempted Zero-Inflated Negative Binomial (ZINB) (ridout2001score, ) for Vul_Count to handle both issues simultaneously.
+However, ZINB models encountered numerical convergence failures with precision loss warnings during maximum likelihood estimation.
+Therefore, we applied NB regression across all outcomes, providing stable parameter estimation while adequately addressing overdispersion.
+5.1.3. High vs. Low Aggregated Scorecard Scores: Comparing Similar Repositories
+To move beyond correlational evidence and establish causality, we employed PSM to simulate experimental conditions by comparing high and low-scoring repositories with similar project characteristics.
+PSM (austin2009balance, ) is a statistical technique that simulates the conditions of a randomized experiment by matching repositories with similar characteristics but differing in their level of security practice adoption.
+Similar project characteristics refer to repositories matched on the same covariates used as control variables in section 3.1.3.
+Repositories are classified as “treated” if their aggregated scores fall within the top quartile, with the remaining projects serving as a control group.
+We tested multiple treatment thresholds (median, tertile, quartile) and selected the top quartile ($\geq$ 3 on a 1-10 scale) to ensure sufficient contrast between treatment and control groups while maintaining adequate sample sizes.
+Using logistic regression on covariates that influence both aggregated scores and security outcomes, we estimated each repository’s propensity score (probability of having a high aggregated score given its characteristics), then matched treated and control repositories with similar propensities.
+We then applied nearest-neighbor matching with a 0.1 standard deviation caliper to pair treated and control repositories with similar propensities.
+We validated matching quality using standardized mean differences (SMD) and calculated Average Treatment Effects (ATE) through difference-in-means with bootstrap confidence intervals (CI) to quantify uncertainty.
+We complement this analysis with NB on post-matched samples (containing only the matched treated and control repositories).
+Table 2. Segment-specific ATEs and percentage impacts. Strongest tercile per variable shown.
+Outcome Controls Segment ATE Impact
+Vul_Count Contributors_CT Many -8.2 28.9%
+Age Mature -11.0 42.2%
+Commit Staleness Fresh -9.4 47.5%
+Downloads High -9.3 35.9%
+Size Large -8.4 27.7%
+Dependency Medium-Dep -9.8 37.0%
+MTTR Contributors_CT Many -350.1 41.7%
+Age Mature -429.8 37.5%
+Commit Staleness Fresh -244.8 37.3%
+Downloads High -269.5 40.5%
+Size Large -315.0 40.2%
+Dependency Few -245.7 28.1%
+MTTU Contributors_CT Many -87.2 19.1%
+Age Mature -86.0 12.8%
+Size Small 146.3 40.0%
+Commit Staleness Stale 212.1 23.5%
+Downloads Few 146.1 38.5%
+Dependency High-Dep -81.0 20.6%
+5.1.4. Impact Across Segments of Each Repository Characteristic
+While our PSM analysis establishes an overall causal effect by controlling for repository characteristics, it estimates one average treatment effect without examining how this effect varies across different repository types.
+For instance, the security benefits of higher aggregated scores might be stronger for large repositories compared to small ones.
+To examine whether this effect differs in magnitude across different levels of repository characteristics, we conducted post-hoc heterogeneity analyses on the matched sample.
+We stratified the matched repositories into terciles based on each control variable using quantile-based binning and calculated segment-specific ATEs, comparing high and low scorecard scores within each segment.
+This approach identifies which repository contexts benefit most from higher aggregated scores.
+We used bootstrap CI to quantify uncertainty around segment-specific ATEs.
 5.2. RQ1: Result
-Figure 3 shows a sharp decline in security outcomes (in Vul_Count, MTTR, and MTTU) as the aggregated Scorecard score increases.
+Table 3. Combined GLM and PSM analysis for RQ1
+Outcome NB Coef NB Coef (PSM) PSM ATE (95% CI)
+Vul_Count -0.20 -0.26 -5.2 [-5.7, -4.7]
+MTTR -0.05 -0.16 -216.8 [-241.8,-193.2]
+MTTU 0.03 -0.04 -52.3 [-60.2, -44.8]
+Data Exploration: Figure 2 shows a sharp decline in security outcomes (indicating improvement) as aggregated Scorecard scores increase.
 The highest bin (7.3–9.1) reflects the top-scoring repositories, as no project achieved a perfect score.
 However, the plot only reflects an overall trend and does not account for confounding factors and associations.
-To formally assess the relationship, we use regression analysis and propensity score matching (PSM).
-Table 3 demonstrates that a higher aggregated Scorecard score is associated with improved security outcomes (negative coefficient).
-In the ZINB model, a coefficient of –1.4 for Vul_Count implies that a unit increase in the aggregated Scorecard score corresponds to a 1.4 count decrease in the expected number of vulnerabilities.
-The finding is also supported by negative ATE estimates (–1.9) in the PSM analysis.
-For MTTU, both the NB model (coefficient –0.2) and PSM estimates (–0.5) indicate that repositories with higher aggregated Scorecard scores update dependencies more rapidly.
-By contrast, the NB model for MTTR shows a downward trend (Figure 3) and a negative coefficient (–0.1).
-However, the PSM analysis yields a positive overall ATE (74.7), indicating that, when controlling for control variables, repositories with higher aggregated Scorecard scores tend to have longer update delays for vulnerable components.
-Table 3. GLMs and PSM analysis between Aggregated Scorecard Score and Security Outcomes
-Dep. Var Model Coef. p-value ATE(Weighted)
-Vul_Count ZINB -1.4 ¡0.001 -1.9
-MTTR NB -0.1 ¡0.001 74.7
-MTTU NB -0.2 ¡0.001 -0.5
-We conducted a segment-based analysis for MTTR to investigate further the discrepancy by stratifying the dataset into subgroups based on repository characteristics using a median split.
-The segment analysis revealed that projects in the upper groups for size, contributors, dependency count, and downloads and the lower group for age tend to exhibit negative ATEs (indicating faster MTTR).
-In contrast, repositories in the lower groups for size, contributors, dependency count, and downloads and in the upper age group show positive ATEs (indicating slower MTTR).
-Overall, the segment analysis reveals that the effectiveness of security practices adoption in reducing MTTR is context-dependent.
-Larger, newer repositories with more contributors, dependencies, and downloads benefit from adopting security practices, resulting in faster MTTR.
-In contrast, smaller but older repositories with fewer contributors, downloads, and dependencies may experience increased overhead from adopting security practices, leading to longer MTTR.
-Summary: Our analysis shows that higher aggregated Scorecard scores are associated with fewer vulnerabilities and shorter MTTU.
-However, while regression models indicate shorter MTTR, causal analysis reveals that repositories with higher aggregated Scorecard scores may take longer to update vulnerable components since project characteristics also influence the relationship.
-Table 4. Distribution of Security Practices Adoption
-Metric % -1 % 0 % 1-10
-Branch Protection 0.005% 85.9% 14%
-Code Review 1.4% 74.6% 24%
-License 0 31% 69%
-Maintained 0.001% 86.7% 13.3%
-Pinned Dependencies 72.2% 0 27.9%
-Table 5. R2, Security practice Importance Ranked and PSM Insights
-Dep. Var R2 for RF Ranked Security Practices PSM Effect (ATE weighted direction)
-Vul_Count 0.38 Code Review, Maintained, Branch Protection, License, Pinned-Dependencies Code Review (+), Others (-)
-MTTR 0.42 Maintained, Pinned-Dependencies, Branch Protection, Code Review, License All (+)
-MTTU 0.2 Maintained, Code Review, Pinned-Dependencies, Branch Protection, Security Policy Code Review (+), Branch Protection(+), Others (-)
-6. RQ2: Associations between security practices and outcomes
+Quantitative Confirmation: Table 3 confirms visual trends (Figure 2) through three analytical approaches: full-sample regression, PSM post-matched sample regression, and PSM causal analysis.
+The columns show full-sample NB coefficients, post-matching NB coefficients, and post-matching ATEs with 95% confidence intervals.
+For Vul_Count, both full-sample (–0.20) and post-matching (–0.26) NB models show negative coefficients, with ATE indicating 5.2 fewer vulnerabilities.
+MTTR demonstrates consistent improvements: full-sample (–0.05), post-matching (–0.16), and ATE (–216.8 days).
+For MTTU, while the full-sample model shows a positive coefficient (0.03), post-matching (–0.04) and ATE (–52.3 days) reveal faster updates, suggesting confounding factors mask the relationship.
+Repository Context-Dependent Effects: While our PSM analysis established overall causal effects, we conducted segment-specific analyses to understand where these effects are strongest.
+Table 2 presents the most pronounced effects within each repository characteristic, with the ”Segment” column identifying the specific subgroup and the ”Impact” column quantifying percentage improvements relative to the control baseline within that segment.
+Across all segments, Vul_Count consistently shows positive security outcomes (i.e., negative ATE values), with improvements ranging from 27.7% to 47.5%.
+The most substantial reductions in open vulnerabilities are observed in mature repositories with fresh commit staleness (47.5% and 42.2%, respectively), indicating that actively maintained and mature repositories with high aggregated scores tend to have fewer vulnerabilities.
+Medium-dependency group repositories show the greatest vulnerability reduction, highlighting the effect of dependency complexity.
+MTTR improvements are also observed across all segments.
+Mature repositories have the largest absolute improvement (429.8 days faster), followed by those with many contributors (350.1 days).
+Repositories with fewer dependencies and fresh commit staleness show meaningful reductions as well (-245.7 and -244.8 days).
+Then, repositories with higher download counts tend to have fewer Vul_Count and faster MTTR.
+MTTU reveals the most heterogeneous pattern, with some segments showing improvements while others exhibit deterioration.
+Small repositories by size and download and repositories with stale commit activity show increases in MTTU, suggesting that top quartile aggregated score repositories that are small, less downloaded, or less active have slower dependency updates compared to low Scorecard score repositories in the same segment category.
+Conversely, repositories with many contributors, mature age, and high dependency counts demonstrate faster updates.
+Summary of RQ1 Findings:
+  • Overall Effects: Higher aggregated Scorecard scores improve security outcomes, with repositories achieving 5.2 fewer vulnerabilities, 216.8 days faster MTTR, and 52.3 days faster MTTU.
+  • Repository Context Matters: The adoption of aggregated security practices improves security outcomes.
+The segment-specific analysis reveals nuanced patterns in how repository characteristics influence security practice effectiveness.
+  • Organizational Capacity: Mature repositories with many contributors consistently show better outcomes, suggesting that well-staffed and mature projects with high adoption of security practices benefit from improved security outcomes.
+  • Project size, activity, and downloads: Project size, recent commit activity, and download volume are associated with improvements in Vul_Count and MTTR.
+However, their absence creates challenges (i.e., smaller, inactive, or less-downloaded projects) and correlates with slower MTTU, highlighting how inactivity and smaller projects with fewer downloads can hinder general updates.
+  • Dependency Complexity: Different dependency levels excel in different areas.
+Medium-dependency projects have the fewest open vulnerabilities, likely balancing manageable complexity.
+Few-dependency projects remediate issues fastest, benefiting from simpler, more transparent dependency trees.
+Large-dependency projects excel at general updates, likely leveraging automation to manage scale effectively.
+While selecting dependencies, organizations should prioritize mature, actively maintained large projects with many contributors, fewer dependencies, and high downloaded packages, as our analysis showed that such projects have fewer vulnerabilities and faster MTTR due to better security performance.
+Conversely, smaller, inactive packages with fewer downloads require enhanced monitoring and proactive security assessments to compensate for slower MTTU.
+Table 4. R2, Security practice Importance Ranked and PSM Insights. Positive ATE indicates increased outcome (worse), negative ATE indicates reduced outcome (better).
+Outcome R2 for RF Ranked Security Practices PSM Effect (ATE direction)
+Vul_Count 0.39 Contributors, Code Review, License, Pinned-Dependencies, Branch Protection, CI-Tests, Dependency-Update-Tool Code Review (+), CI-Tests (mixed), Branch Protection (+), Dependency-Update-Tool (+), Others (-)
+MTTR 0.50 Contributors, Code Review, License, CI-Tests, Pinned-Dependencies Code Review (+), Others (-)
+MTTU 0.32 Contributors, License, Code Review, CI-Tests, Pinned-Dependencies Code Review(+), CI-Tests (mixed), Others (-)
+6. RQ2: Impact of individual security practices and outcomes
 In this section, we discussed our RQ2: Which security practices are most strongly associated with security outcomes?
-Section 6.1 discusses our methodology to answer our RQ2, and Section 6.2 discusses our result.
+Section 6.1 discusses the methodology for RQ2, and Section 6.2 discusses the result.
 6.1. RQ2: Methodology
 For RQ2, we investigate the association between individual security practices and security outcomes.
-Unlike RQ1, where we analyzed aggregated Scorecard scores as a predictor, for RQ2, we analyzed twelve distinct security practices as predictors.
-The large number of predictors, non-normal distribution, and non-linear relationships among variables (section 4) introduce additional complexity to the dataset.
-We selected a tree-based regression model as these models are well-fitted for datasets with complex, non-linear relationships and high-dimensional feature interactions.
+Unlike RQ1, where we analyzed aggregated Scorecard scores as a predictor, for RQ2, we analyzed eleven security practices as predictors after exclusion (Section 3.2.3 and 4.2) alongside six control variables.
+Due to the large number of predictors, non-normal distribution, and non-linear relationships (Section 4), we selected a tree-based regression model suited for handling complex, high-dimensional data.
 We also conducted the feature importance ranked analysis to identify the security practices that most strongly contribute to predicting security outcomes.
@@
 6.1.1. Regression Model:
-Among tree-based models, we chose Random Forest (RF) (dewi2019random, ) due to its ensemble learning approach, which aggregates multiple decision trees to improve model performance and reduce overfitting.
-Unlike single decision trees, which can be highly sensitive to variations in training data, RF mitigates the issue by averaging predictions across multiple trees.
-We also utilized Gini Importance (Mean Decrease in Impurity) as a feature importance technique to quantify each security practice’s contribution toward reducing prediction error.
-Gini Importance measures a predictor’s contribution by assessing how often the predictor is used for splitting the data in the decision trees of the RF model, highlighting the practice’s contribution in accurately predicting security outcomes.
-We first preprocessed our data, converting categorical security practices into appropriate numerical or binary formats (section 4.1).
-We then split the dataset into training (70%) and testing (30%) subsets.
-The model was trained on the training dataset using default hyperparameters.
-Five-fold cross-validation was applied to estimate and reduce bias, calculating mean squared error (MSE) and R² scores.
-The final model’s performance and feature importance scores were validated on the reserved testing set (Table 5).
-6.1.2. Causal Analysis
-We followed the methodology outlined in Section 5.1.2 for our causal analysis.
-We conducted the PSM analysis only on the top 5 security practices ranked by the Gini feature importance measure.
-The top 5 was selected because, after that, the feature importance value was almost near 0, indicating those practices did not have a strong association with security outcomes.
-The aggregated Scorecard score is computed as a weighted average of 18 individual security practices; only 25 repositories resulted in a score of zero.
-On the contrary, individual security practices exhibit a higher rate of non-adoption or security misconfiguration (score -1, score 0), as shown in Table 4 (top 5 based on feature ranked).
-The disparity occurs because aggregation combines multiple practices, ensuring a nonzero score even when some practices are not adopted.
-In contrast, an individual practice remains at 0 or -1 if the practice is not implemented.
-As a result, individual security practices had a higher proportion of repositories with a score of 0 or -1, leading to a highly imbalanced dataset for PSM analysis.
-To mitigate imbalance and ensure balanced treatment and control groups, we applied a sampling approach for the top 5 security practices.
-We first included all repositories with scores between 1 and 10, then randomly sampled an equal number of repositories with a practice score of zero to create a representative baseline.
-We then created a binary treatment variable by splitting the combined dataset based on an adaptive threshold (below and above the median score).
-However, for Maintained and License metrics, the median score was equal to the maximum observed score (indicating minimal variance), and we adjusted the threshold to the midpoint between the minimum and maximum scores to ensure separation between high and low adoption groups.
-After constructing balanced groups, we followed the technique discussed in Section 5.1.2, where we used nearest-neighbor matching with a caliper adjustment and computed the ATE score (ATE weighted).
+We employed Random Forest (RF) regression due to its robustness to non-linear relationships and ability to handle high-dimensional data without overfitting.
+Following preprocessing (section 4.1), the dataset was split into training (70%) and testing (30%) subsets using random sampling.
+The model was trained on the training dataset using hyperparameters (n_estimators=100, random_state=42).
+To ensure robust performance estimation and reduce bias, five-fold cross-validation was applied exclusively to the training data, calculating R² scores(Table 4).
+6.1.2. Security Practice Ranking
+: Feature importance was quantified using Gini Importance, extracted directly from the trained Random Forest model’s feature importances attribute, which measures each feature’s contribution to reducing node impurity across all decision trees in the ensemble.
+Features were ranked by their grouped importance scores using dense ranking methodology (method= dense) within each dependent variable.
+Since the preprocessing step decomposed high-missing features (CI-Tests, Pinned-Dependencies) into separate components, we grouped all feature importance scores by their base security practice to provide more interpretable ranking results.
+The method involved summing the decomposed components for CI-Tests and Pinned-Dependencies, while non-decomposed features retained their individual importance scores.
+Using the ranked importance scores of 11 practice metrics, we selected the top-performing security practices that collectively accounted for approximately 80% of the cumulative feature importance across all three outcome metrics (ucla2023efa, ).
+The approach provides interpretable insights into which security practices have the strongest overall predictive power for each outcome variable (Table 4).
+6.1.3. High vs. Low Security Practice Scores: Comparing Similar Repositories
+We conducted PSM analysis on the top-ranked security practices from Table  4.
+We focused on these top practices because the remaining ones demonstrated negligible importance values while the selected top features collectively accounted for approximately 80% of the cumulative feature importance.
+However, individual security practices exhibit extreme binary distributions, with most repositories scoring either 0 (not adopted) or 10 (fully implemented), creating an extreme class imbalance that violates PSM overlap assumptions.
+Unlike aggregated Scorecard scores, which combine multiple practices to ensure nonzero values, individual practices remain at their adoption extremes.
+To enable valid causal inference, we applied balanced sampling to create sufficient propensity score overlap.
+Our PSM methodology employed a universal threshold of 5.0, classifying repositories with scores above this threshold as treated units and those at or below as control units.
+Following balanced sampling, after constructing balanced groups, we followed the technique discussed in Section 5.1.3, where we used nearest-neighbor matching with a caliper adjustment and computed ATE.
+Unlike RQ1, where segmentation revealed how repository characteristics impact outcome, we could not perform a similar analysis in RQ2 due to data imbalance (Section 6.1).
+The limited sample size in the PSM analysis further constrained our ability to isolate project-specific influences on security outcome metrics.
 6.2. RQ2: Result
-Our analysis employed RF regression, complemented by feature importance analysis and Propensity Score Matching (PSM), to identify security practices most strongly associated with security outcomes.
-Table  5 summarizes our findings, illustrating the variance explained (R²) by security practices and each practice’s directional effects based on the PSM analysis’s ATE-weighted values.
-The R² value indicates that security practices collectively explain a moderate portion of the variance in Vul_Count (R² = 0.38) and MTTR (R² = 0.41), and a relatively smaller proportion for MTTU (R² = 0.2).
-Notably, the lower R² for MTTU implies that additional external factors or project-specific characteristics beyond those captured in our current model likely contribute to MTTU.
-From the feature importance analysis, Code Review, Maintained, Pinned Dependencies, and Branch Protection practices consistently rank highly across all three outcome metrics.
-License considerations showed relevance for both Vul_Count and MTTR, whereas Security Policy appears influential for MTTU, highlighting the variability of practice impacts across distinct outcome measures.
-In the PSM analysis of individual security practices, we did not consistently observe a negative (i.e., improving) effect across all security outcomes.
-Most security practices show a negative ATE for Vul_Count and MTTU (Table  5), implying fewer vulnerabilities and faster update frequency.
-However, Code Review consistently exhibits a positive relationship with all three security outcomes, suggesting that the practice aids in detecting more vulnerabilities.
-As emphasized in CISA’s Secure by Design guidance (cisa2023securebydesign, ), a rise in reported vulnerabilities can reflect a healthy security posture due to code review and testing by the community.
-For MTTR, all practices exhibit a positive ATE, indicating longer remediation times for repositories adopting these practices.
-Figure 3 shows that 90% npm repositories fall into low-security score bins (0.0–3.6), where MTTR is high (1200 days), suggesting these projects do not prioritize dependency security or vulnerability remediation.
-Table 4 further highlights that many npm packages lack security practices adoption, reinforcing this trend.
-While higher aggregated security scores and individual practices are associated with fewer Vul_Count and shorter MTTU, MTTR remains positive for both cases, suggesting MTTR is influenced by factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity.
-Unlike RQ1, where segmentation revealed how repository characteristics impact MTTR, we could not perform a similar analysis in RQ2 due to data imbalance (Section 6.1).
-The limited sample size in the PSM analysis further constrained our ability to isolate project-specific influences on MTTR.
-Therefore, we cannot conclusively confirm whether the observed increase in MTTR consistently indicates slower remediation or whether the effect might vary depending on project-specific characteristics, maintainability, or dependency complexity.
-Overall, Vul_Count , MTTR, and MTTU exhibit similar trends as security outcomes metrics.
-However, MTTR is more complex and may have an influence on external factors such as project maintainability, resources, and dependency complexity that need further research.
-Summary: Our analysis shows that Code Review, Maintained, Pinned Dependencies, and Branch Protection are strongly associated with security outcomes.
-While individual security practice is associated with fewer vulnerabilities and faster dependency updates, MTTR remains high, suggesting that factors beyond security practice adoption, such as project maintainability, resources, and dependency complexity, influence MTTR.
+Quantitative Confirmation: Table  4 summarizes our RQ2 findings, analyzing individual security practices’ relationships with security outcomes.
+Security practices collectively explain a moderate portion of variance across outcomes: R² = 0.39 for Vul_Count , R² = 0.50 for MTTR, and R² = 0.32 for MTTU.
+The higher explanatory power for MTTR suggests stronger relationships between security practices and MTTR compared to Vul_Count or MTTU.
+Important Security Practices: The analysis reveals consistent patterns across outcome metrics.
+Contributors emerge as the most influential factor across all three outcome metrics, followed by Code Review and License practices.
+CI-Tests and Pinned-Dependencies consistently rank among the top practices.
+The ranking pattern suggests that collaborative human factors (Contributors) and development workflow practices (Code Review, CI-Tests) have an influential, measurable impact on security outcomes.
+Branch Protection and Dependency Update Tool ranks among the top seven features required to reach the 80% cumulative importance threshold for Vul_Count.
+In contrast, MTTR and MTTU reach the threshold with five features, but including these two additional practices raises the cumulative importance to nearly 92% for MTTR and MTTU, indicating these practices also have a stronger impact on the outcomes.
+Impact of Security Practice: The PSM analysis reveals complex and counterintuitive relationships between individual security practices and outcomes.
+Code Review consistently exhibits a positive ATE across all three security outcomes, indicating associations with more vulnerabilities, slower MTTR, and MTTU.
+This paradoxical finding aligns with CISA’s Secure by Design guidance (cisa2023securebydesign, ), a rise in reported vulnerabilities can reflect a healthy security posture due to code review and testing by the community.
+CI-Tests show a mixed pattern: projects with pull-request (PR) infrastructure tend to have faster MTTR but also higher vulnerability counts and slower MTTU, possibly due to added complexity from collaboration.
+However, projects with PR infrastructure that also achieve high CI-Test scores consistently perform better across all security metrics.
+The finding suggests that PR workflows alone are insufficient; security benefits arise when effective CI practices are fully implemented.
+Beyond Code Review and CI-Tests, other practices show more intuitive associations.
+Contributors, License, and Pinned Dependencies are associated with negative ATEs, suggesting improved outcomes if adopted by repositories.
+For Vul_Count, the Branch Protection and Dependency Update Tool shows positive ATEs, suggesting these practices enhance vulnerability detection capabilities.
+Our analysis shows that Contributors, Code Review, License, CI-Tests, and Pinned-Dependencies are the most influential security practices across all outcome metrics.
+While adoption of Code Review practice shows positive associations with more vulnerabilities and slower MTTR and MTTU, Contributors, License, and Pinned Dependencies demonstrate negative ATEs, indicating improved security outcomes.
+CI-Tests show mixed effects depending on CI-test adoption.
 7. Discussion
-The section discusses the novelty of our study (section 7.1) and the implications of our findings (section 7.2) and explores the threat to the validity (section 7.3) of our study.
-7.1. Novelity of our study
+The section discusses the research implications (section 7.1) and explores the threat to the validity (section 7.2) of our study.
+7.1. Research Implication
+Defining and validating security outcome metrics is challenging; our work represents a step toward advancing that goal.
+Our study provides actionable insights for practitioners, policymakers, and researchers by investigating measurable security outcome metrics to assess the effectiveness of security practices.
+Holistic security practice adoption:.
+RQ1 demonstrates that aggregated scores consistently improve security outcomes, while RQ2 reveals context-dependent and sometimes counterintuitive effects of individual practices.
+Our findings suggest that comprehensive security practice adoptions are more effective than isolated measures, as coordinated practices generate synergistic effects that individual actions alone cannot achieve.
+Context-Aware Adoption: The impact of security practice on security outcomes varies by project characteristics, requiring tailored strategies.
+While selecting dependencies, organizations should prioritize mature, actively maintained large projects with large contributors, fewer dependencies, and high downloaded packages, as our analysis showed that projects have fewer vulnerabilities and faster MTTR due to better security performance.
+Conversely, smaller or inactive packages with fewer downloads require enhanced monitoring and proactive security assessments to compensate for slower MTTU.
+Strategic Practice Prioritization: Despite individual variability, Code Review, CI-Tests, License, Contributors, and Pinned Dependencies consistently demonstrate high feature importance across security outcomes.
+Our findings suggest that practitioners can adopt these empirically validated practices within their workflow, while recognizing that practices like Code Review, CI-test positive association reflect enhanced detection capabilities rather than security degradation.
+To ensure that security practice adoption is yielding tangible benefits, teams should regularly monitor outcome metrics.
+Continuous tracking of these indicators enables practitioners to assess the effectiveness of implemented practices and iteratively refine their security strategies based on empirical feedback.
+Evidence-Based Policy Development: The quantified relationships between security practices and security outcomes provide empirical evidence for security frameworks and regulations.
+Policymakers can leverage these findings to develop guidelines that promote frameworks emphasizing coordinated, measurable practice adoption rather than checklist-based compliance.
+By aligning regulatory efforts with empirically supported metrics, policy can more effectively incentivize secure development behaviors across the software ecosystem.
+Future Work: Our findings highlight the need to better understand the temporal feedback loop between security practices and outcomes.
+As projects adopt more practices, they may initially report more vulnerabilities and delays in dependency updates and fixing vulnerabilities due to improved detection rather than increased risk.
+The complexity calls for longitudinal studies that track how individual and aggregated practices affect security outcome metrics.
+For researchers, future work should model these dynamics in the context of project maturity, contributor activity, and development cycles to clarify whether observed security trends reflect degradation or deliberate, ongoing improvement.
+A comprehensive understanding of security outcome metrics is an open research challenge; therefore, researchers should explore new security outcome metrics beyond those used in this study.
+7.2. Threat to validity
+Our regression and causal analyses provide evidence supporting associations and causality between aggregated Scorecard scores, security practices, and security outcomes within our dataset.
+However, several inherent biases in our dataset must be acknowledged.
+First, our analyses are based on snapshot data capturing repository Scorecard metric scores and characteristics at a single point in time, lacking temporal granularity.
+Without longitudinal or time-series analysis, uncertainty remains about whether changes in security practices precede or directly influence observed security outcomes, thereby limiting the generalizability of our findings beyond our dataset.
+Second, selection bias may be present, as repositories adopting security practices could differ in unmeasured factors such as organizational policies, economic incentives, developer motivations, or project complexity.
+Although PSM mitigates confounding through observed covariates, unobserved variables may still influence both practices and outcomes.
+Apart from that, our dataset exhibits other imbalances, as repositories have security practice scores of either 0, or -1, reflecting the limited adoption of these practices across the npm ecosystem.
+The real-world imbalance, characterized by the few repositories that adopt security practices, may introduce bias into our models, potentially skewing predictions or causal estimations.
+However, this imbalance highlights a widespread lack of security practice adoption among npm repositories.
+Hence, the imbalance reflects the actual state of adoption rather than a methodological flaw, highlighting the need for targeted interventions to increase the adoption of security practices among npm repositories.
+At the same time, the metric collection also needs improvement, a team may be following a practice, but evidence of that practice may not be captured in the automated metrics.
+8. Related Work
+This section highlights prior related studies, existing security guidelines, frameworks, and security outcome metrics.
 Prior studies (zahan2023software, ; sonatype_8, ) have utilized OpenSSF Scorecard metrics alongside vulnerability count data to understand the relationship between security practices and security outcome metrics.
-Building on this foundation, we expand security outcome metrics to include MTTR and MTTU.
-Our findings showed that Vul_Count, MTTR, and MTTU are associated with security practices.
-However, MTTR is more complex and may have an influence on external factors such as project maintainability, resources, and dependency complexity.
-Additionally, our study further advances existing research by addressing methodological limitations in prior work.
 Zahan et al.
-(zahan2023software, ) constructed a small dataset of 2,422 PyPI and npm packages and analyzed the relationship between OpenSSF Scorecard metrics and vulnerability counts.
-The authors’ findings revealed a positive association between security practice adoption and vulnerability counts, alongside a low explanatory power ($R^{2}$ 12%).
-The authors attributed these results to dataset size limitations, lack of vulnerable data, and the absence of control variables related to project characteristics.
-In contrast, our study leverages a larger and more representative dataset, encompassing approximately 146,951 npm package repositories, which mitigates biases introduced by small sample sizes and enhances generalizability.
-Additionally, we incorporate control variables, including project age, size, contributors, and dependency count and downloads, to account for potential confounding factors influencing security outcomes.
-Furthermore, Zahan et al.’s dataset had a temporal mismatch between security practice adoption and vulnerability data.
-The authors collected security practice scores at a single point in time but included all historically reported vulnerabilities throughout a package’s lifetime, potentially distorting the actual relationship between security practices and security outcomes.
-Our methodology addresses this inconsistency by aligning security practice adoption with vulnerabilities that actively existed at the time of our study.
-By ensuring a concurrent snapshot of security practices and vulnerabilities, our approach provides the temporally aligned evaluation of security practice effectiveness.
-Additionally, we addressed the issue of data consistency in study (zahan2023software, ), stemming from repositories containing multiple packages, each potentially contributing different Vul_Count  (Section  3.2), despite sharing identical Scorecard scores.
-Moreover, vulnerabilities in prior studies (zahan2023software, ) did not include vulnerabilities from dependencies or entire repositories but rather only individual packages, introducing further inconsistencies.
-Such discrepancies result in duplicate predictors with varying outcome values, potentially biasing predictive models.
-To mitigate these issues, we filtered our dataset to include only repositories linked uniquely to a single package.
-We considered vulnerabilities explicitly associated with the repository itself and those inherited from dependencies.
-In another study, Sonatype (sonatype_8, ) employed a random forest classification model to classify repositories as vulnerable or non-vulnerable using Security Scorecard metrics and project attributes.
-However, a binary classification approach is inherently limited in capturing relationships between security practices and security outcomes.
-The classification does not effectively measure the magnitude or direction of associations, nor does it indicate the relative strength or complexity of individual predictors.
-In contrast, our study utilizes regression and causal analysis, methods specifically designed to explore and quantify associations, thus providing insights into how variations in security practices systematically relate to security outcomes.
-Summary: Overall, our methodological refinements, including extensive dataset size, the inclusion of project characteristics, causal analysis, targeted dataset inclusion strategies, and the introduction of MTTR and MTTU as additional novel security outcome metrics, offer a more precise and contextually generalizable evaluation of the relationship between security practices and outcomes.
-7.2. Implications of our study
-Our study provides actionable insights for practitioners, policymakers, and researchers by establishing measurable security outcome metrics to assess the effectiveness of security practices.
-For practitioners, our findings highlight that Code Review, Maintained, Pinned Dependencies, and Branch Protection are strongly associated with security outcomes.
-Organizations can prioritize these practices to improve their project security.
-By periodically tracking security outcome metrics, practitioners can evaluate whether their security interventions are effectively reducing vulnerabilities and improving remediation efficiency.
-Our research supports policymakers by providing empirical evidence on the effectiveness of security practices, informing data-driven policy decisions, and enhancing regulatory frameworks.
-For researchers, our study underscores the importance of incorporating longitudinal data, expanded covariate sets, and more granular temporal analyses to understand better the dynamic interactions between security practices, external factors, and security outcomes.
-Additionally, future research should explore new security outcome metrics beyond vulnerability counts and remediation times to develop a more comprehensive framework for measuring security improvements in software development.
-7.3. Threats to validity
-Our regression and causal analyses provide evidence supporting associations between aggregated Scorecard scores, security practices, and security outcomes within our dataset.
-However, several inherent biases in our dataset must be acknowledged.
-First, our analyses are based on snapshot data capturing repository Scorecard metrics score and characteristics at a single point in time, lacking temporal granularity.
-Without longitudinal or time-series analysis, uncertainty remains about whether changes in security practices precede or directly influence observed security outcomes, thereby limiting generalizability beyond our dataset.
-Second, selection bias may be present, as repositories that adopt higher security practices could systematically differ from those with lower adoption in unmeasured aspects such as organizational resources, developer expertise, or security culture.
-The unobserved factors may confound the relationships we examine, leading to potential over or underestimations of the true impact of security practices.
-Third, although PSM attempts to mitigate confounding factors by simulating randomized experiments, PSM inherently relies on observed covariates.
-Therefore, unobserved or missed variables such as organizational policies, economic incentives, developer motivations, or project complexity could influence both security practices and security outcomes.
-Additionally, the consistently observed positive ATE for MTTR suggests that security practices may introduce operational overhead, and the relationship is context-dependent, where repositories with project characteristics may experience varying impacts.
-Apart from that, our dataset exhibits other imbalances, as most repositories have security practice scores of either 0 or -1, reflecting the limited adoption of these practices across the npm ecosystem.
-The real-world imbalance, characterized by few repositories adopting security practices, may introduce bias into our models, potentially skewing predictions or causal estimations.
-However, such imbalance accurately mirrors the current state of practice adoption within the npm, highlighting a widespread lack of security practice adoption among npm repositories.
-Hence, the imbalance reflects the actual state of adoption rather than a methodological flaw, highlighting the need for targeted interventions to increase security practice adoption among npm repositories.
-Finally, while our filtering approach improves data consistency by retaining only repositories uniquely linked to a single npm package and aligning vulnerability data at the repository level and MTTR and MTTU data at the package level, some inconsistencies may persist.
-For example, repositories that maintain multiple packages from different ecosystems or use different languages may still introduce variations in MTTR and MTTU value, potentially affecting model accuracy.
-Summary: Overall, while our study provides insights into the relationships between security practices and security outcomes, the limitations highlight the need for further research.
-Future studies incorporating longitudinal data, expanded covariate sets, and more granular temporal analyses will be essential to better capture the dynamic interactions between security practices, external factors, and security outcomes.
-8. Related Works
-This section highlights prior related studies, existing security guidelines, frameworks, and security outcome metrics.
-Security Practice Frameworks: Executive Order (EO) 14028 emphasized adopting and self-attesting security practices to improve software security.
+(zahan2023software, ) analyzed 2,422 PyPI and npm packages and found a positive association between Scorecard metrics and vulnerability counts at the package level, but with low explanatory power ($R^{2}$ $<$ 12%).
+They attributed this to limited dataset size, sparse vulnerability data at the package level, and missing control variables.
+Sonatype (sonatype_8, ) used a random forest classifier to predict vulnerable repositories based on Scorecard metrics and project attributes.
+However, binary classification limits understanding of the strength and direction of relationships between practices and outcomes.
+In contrast, our study employs regression and causal analysis to quantify associations, uses a larger, more representative dataset of 145,817 npm packages, incorporates project-level outcome metrics, and controls for key project characteristics to improve validity and generalizability.
+Security Practice Frameworks: Executive Order (EO) 14028 emphasized adopting security practices to improve software security.
 To align with the EO, government agencies and industry experts have introduced various software supply chain security frameworks to enhance software supply chain security.
@@
 Beres et al.
-(beres2009using, ) studied the window of exposure’ (vulnerability metric), which is the length of interval of time needed by the security team to remediate a vulnerability once it has been disclosed.
+(beres2009using, ) studied the window of exposure (vulnerability metric), which is the length of the interval of time needed by the security team to remediate a vulnerability once it has been disclosed.
 However, the scarcity of reported vulnerabilities, compounded by silently fixed issues, limits their reliability in assessing security outcomes (alhazmi2007measuring, ; dunlap2023finding, ).
@@
 9. Conclusion
-In this study, we investigated how aggregate Scorecard score and 12 Scorecard security practices metrics impact security outcome metrics.
-Our regression model and causal analysis reveal that higher aggregated Scorecard scores are associated with fewer vulnerabilities and shorter MTTU.
-However, while regression models indicate shorter MTTR, causal analysis suggests that factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity, influence MTTR.
-Feature ranking highlights Code Review, Maintained status, Pinned Dependencies, and Branch Protection as top security practices associated with security outcomes, though their associations with security outcomes vary in direction.
+In this study, we investigated how the aggregate Scorecard score and Scorecard security practices metrics impact security outcome metrics.
+Our regression model and causal analysis reveal that higher aggregated Scorecard scores are associated with fewer vulnerabilities, shorter MTTR, and shorter MTTU.
+However, while regression models and causal analysis indicate better outcomes, our segment analysis to understand the impact of project characteristics suggests that factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity, influence security outcomes.
+Feature ranking highlights Code Review, Contributors, License, CI-Tests, and Pinned Dependencies as top security practices associated with security outcomes, though their associations with security outcomes vary in direction.
 Additionally, MTTU exhibits trends similar to vulnerability count and MTTR, making it a potential proxy metric when vulnerability and MTTR data are unavailable.
-Our research aligns with federal cybersecurity initiatives (EO 14144), by providing empirical evidence to guide security policy and industry practices.
-Future work should incorporate longitudinal data and additional security outcome metrics to refine further our understanding of how security practices impact software security over time.